### v1

# Perspective on Moreau–Yosida Regularization in Density-Functional Theory

Abstract
Within density-functional theory, Moreau–Yosida regularization enables both a reformulation of the theory and a mathematically well-defined definition of the Kohn–Sham approach. It is further employed in density–potential inversion schemes and, through the choice of topology for the density and potential space, can be directly linked to classical field theories. This perspective collects various appearances of the regularization technique within density-functional theory alongside possibilities for their future development.
I Introduction
It has been more than 10 years since Moreau–Yosida (MY) regularization was introduced into density-functional theory (DFT) [1], and it has arguably turned out to be an unexpectedly fruitful extension. First introduced as a conceptual tool to fix differentiability issues of the exact functional and to make the exchange–correlation potential mathematically well defined, it was later realized that it can be seen as the theoretical basis of density–potential inversion techniques. In order to connect it to existing methods, like the Zhao–Morrison–Parr inversion scheme [2], the topology of the density and potential spaces needed to be adjusted in a way that is critically different from the standard setting of DFT introduced by Lieb [3] based on Lebesgue spaces $L^{p}$. This recent realization [4] marked an important shift of perspective, since it became apparent that the topology of those spaces can entail physical meaning. More concretely, the duality map that canonically connects the density space and the potential space as its dual, is just the Poisson equation of electrostatics if the spaces are chosen as special Sobolev spaces. In this way, new physical meaning and connections to classical fields enter the theory and need to be interpreted accordingly. At this stage of development, many exciting routes can be taken, be it about the precise choice of topology, the extension of the spaces to include vector potentials, or the utilization of the regularized theory for computational purposes.
This perspective article aims at showing the current state of the field as well as mapping out several directions for continued research that we deem promising. We believe that, as the theory matures, it can establish itself as a mathematically rigorous but also computationally useful extension of DFT. Ideally, this means offering new and enhanced approximations for electronic-structure calculations, while also covering a wider range of physical phenomena. These and many more instances of MY regularization in DFT are discussed in this perspective article, where we also make the relevant connections to the literature.
Since the original intention of introducing MY regularization into DFT was to make the theory mathematically sound, it is necessary to introduce a certain amount of mathematical concepts. We try to present this in the preliminary Section II with minimal requirements, but the interested reader might want to follow up on certain details by consulting the vast mathematical literature on the fields that we mainly rely on. These are functional and convex analysis formulated on Banach and Hilbert spaces, as well as the more specialized theory of Sobolev spaces, and the reference list includes our favorite textbooks [5; 6; 7; 8; 9; 10; 11]. After covering the most important mathematical topics in Sections II.1, II.2, and II.3, we continue in Section II.4 with a formalized presentation of DFT that introduces the universal density functional for a fully interacting system and an auxiliary, non-interacting system. This is complemented by Section II.5 on the Kohn–Sham (KS) method that we introduce with an additional offset potential to capture a priori information such as external binding potentials. Section III is concerned with introducing MY regularization into DFT in three different ways—namely, based on the universal density functional, based on the total energy functional, and based on density–potential mixing. The next two sections describe the main use cases for MY regularization in DFT to date, a technique for density–potential inversion in Section IV and a rigorous formulation of the KS method that takes a central role in the application of DFT in Section V. Section VI shows how certain formulations of DFT, such as Hartree theory and Maxwell–Schrödinger DFT, get automatically regularized through the inclusion of a mean-field energy contribution. Section VII deals with the numerical realization of density–potential inversion in the KS setting for periodic systems. Finally, we conclude with an extended outlook in Section VIII.
II Preliminaries
II.1 Functional analysis preliminaries
Our development of MY regularization applied to DFT relies heavily on topological properties of the involved density and potential spaces, the Banach space $X$ and its topological dual $X^{*}$. We keep them as abstract spaces for most of our presentation, since this allows the application of the results to a huge variety of different settings. It will be indispensable to repeat important results from functional analysis, especially from the theory of Banach spaces, since this will be the first pillar on which we base DFT mathematically. The properties of the spaces $X$ and $X^{*}$ are collected in three different assumption sets of increasing restrictiveness to which later results will refer.
The basic setting for the whole presentation is a real Banach space $X$ with the norm denoted as $\|\cdot\|_{X}$. Banach spaces have been developed from the beginning as spaces of functions [12] and as such typically are infinite dimensional, which complicates their mathematical study. The feature that makes a normed vector space a Banach space is that it is complete, meaning that every Cauchy sequence in $X$ has a well-defined limit within the space $X$. Convergence to a limit value usually means strong convergence—that is, convergence in norm $\|\cdot\|_{X}$, where $x_{n}\to x$ in $X$ if and only if
$\lim_{n\to\infty}\|x_{n}-x\|_{X}=0.$ (1)
Since the norm of the space naturally induces a topology (a notion of ‘closeness’ and ‘continuity’), we are able to define the topological dual of $X$ as the space of linear and continuous maps from $X$ to $\mathbb{R}$ (called ‘functionals’),
$X^{*}=\{u:X\to\mathbb{R}\mid u\;\text{linear and continuous}\}.$ (2)
Now, $X^{*}$ gets equipped with a norm itself, the operator norm on $X$, that we will accordingly denote with $\|\cdot\|_{X^{*}}$,
$\|u\|_{X^{*}}=\sup_{\|x\|_{X}\leq 1}|u(x)|.$ (3)
The completeness of $X$ then carries over to $X^{*}$, making it a Banach space as well. With the help of $X^{*}$ we can define weak convergence $x_{n}\rightharpoonup x$ in $X$ as the property that, for all $u\in X^{*}$, it holds that
$\lim_{n\to\infty}u(x_{n}-x)=\lim_{n\to\infty}u(x_{n})-u(x)=0.$ (4)
As suggested by the terminology, strong convergence of a sequence implies weak convergence. Further, let
$J(x)=\{u\in X^{*}\mid u(x)=\|x\|_{X}^{2}=\|u\|_{X^{*}}^{2}\}$ (5)
be a set-valued map that maps each element in $X$ to a subset of $X^{*}$, called the duality map. The Hahn–Banach theorem guarantees that the image of every $x\in X$ is non-empty and it is easy to see that $J$ is homogeneous (i.e., $J(\alpha x)=\alpha J(x)$ for every $\alpha\in\mathbb{R}$), but we would like to impose further topological properties on $X$ so as to make the duality mapping nicer. A Banach space for which $J(x)$ is always single-valued (and thus has not to be considered as a set-valued map) is called smooth. In a smooth Banach space, the duality map is unique and given by the Gâteaux (functional) derivative of the convex function $\phi(x)=\frac{1}{2}\|x\|_{X}^{2}$,
$J(x)=\nabla\phi(x).$ (6)
Interestingly, the smoothness of a space $X$ implies that its dual $X^{*}$ is strictly convex (the line segment between any two points $u\neq v$, $\|u\|_{X^{*}}=\|v\|_{X^{*}}=1$, lies strictly within the unit ball). This is an expression of the general connection of smoothness properties of a Banach space to convexity features on the dual side. The next important property for our space $X$ is that of reflexivity, which means that the bidual $X^{**}=(X^{*})^{*}$ is isometrically isomorphic to the space $X$ itself and can be identified with it. Hence, if $X$ is reflexive, then $X$ can be recovered from $X^{\ast}$ as $X=X^{**}=(X^{*})^{*}$.
If $X$ is reflexive, then $X^{*}$ is also reflexive. We may therefore set up a duality map $J^{\ast}$ from $X^{\ast}$ to $X=X^{\ast\ast}$. For every $u\in X^{*}$, we can consider an $x\in J^{*}(u)=\{x\in X\mid u(x)=\|x\|_{X}^{2}=\|u\|_{X^{*}}^{2}\}$ and with this recover $u\in J(x)$, so $J$ is also surjective if $X$ is reflexive. For a surjective function we can consider its set-valued inverse and we thus get $J^{-1}=J^{*}$. Now note that for $X^{*}$ also smooth ($X$ strictly convex) we have $J^{-1}$ single-valued as well, so $J$ is indeed bijective. We collect all this properties for the minimal assumptions for our Banach spaces $X$ and $X^{*}$:
Assumption 1.
The real Banach space $X$ is reflexive, with $X$ and $X^{*}$ strictly convex. The corresponding duality map $J:X\to X^{*}$ is single-valued with a well-defined inverse $J^{-1}:X^{*}\to X$.
The prototypical example for such Banach spaces, which is also relevant in the context of DFT, are the Lebesgue spaces $L^{p}$, $1<p<\infty$. It is important to note that $p=1$ and $p=\infty$ must be excluded to fulfill Assumption 1, since those spaces are neither reflexive nor strictly convex. At this point, it already becomes obvious that we need to adapt the typical choice of density space in DFT [3], $L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$, since by including $L^{1}$ the space becomes non-reflexive and thus cannot fulfill Assumption 1. It is precisely for this reason that, for example, Eschrig [13] in his discussion of thermal DFT switches to the reflexive space $L^{3}(\Omega)$ on a bounded three-dimensional periodic domain, which anyway entails $L^{3}(\Omega)\subset L^{1}(\Omega)$.
Assumption 1 has been our standard assumption in several previous publications on Moreau–Yosida regularization in DFT [14; 15; 4] and is sufficient for many use cases. Yet, in some instances, additional regularity is required and we add the condition that $X^{*}$ is also uniformly convex. Note that uniform convexity implies reflexivity. A space $X^{*}$ is called uniformly convex if, for each $\varepsilon\in(0,2]$, there is a $\delta(\varepsilon)>0$ such that, for all $u,v\in X^{*}$ with $\|u\|_{X^{*}}=\|v\|_{X^{*}}=1$ and $\|u-v\|_{X^{*}}=\varepsilon$, we have $1-\frac{1}{2}\|u+v\|_{X^{*}}\geq\delta(\varepsilon)$. Adding the requirement of uniform convexity, we obtain:
Assumption 2.
$X$ and $X^{*}$ fulfill all requirements of Assumption 1 and additionally $X^{*}$ is assumed uniformly convex. The duality map $J:X\to X^{*}$ is then single-valued and uniformly continuous.
Finally, we can tighten those assumptions even more and move to a Hilbert space $X$, equipped with an inner product denoted as $\langle\cdot,\cdot\rangle_{X}$. Hilbert spaces are automatically uniformly convex and thus also reflexive. Since then also $X^{*}$ is a Hilbert space, we have all the properties in Assumption 2 fulfilled. For Hilbert spaces, the duality map $J:X\to X^{*}$ is exactly the isometric isomorphism from the Riesz representation theorem and linear as such. Nevertheless, we do not want to identify a Hilbert space with its dual in the given context. The original work introducing Moreau–Yosida regularization to DFT of Kvaal et al. [1] considered a Hilbert-space setting and it is also employed elsewhere [16; 17; 18; 19]. We also highlight a comprehensive chapter by Kvaal [20] in a recent textbook on DFT.
Assumption 3.
The Banach space $X$ and consequently also $X^{*}$ are Hilbert spaces. The duality map $J:X\to X^{*}$ is then linear and Lipschitz continuous with constant 1. Even though $X=X^{*}$, we will not assume $J$ chosen to identify the space with its dual.
For each $u\in X^{*}$, the element $x=J^{-1}(u)\in X$ is then just the unique element given from the Riesz representation theorem and we have for every $y\in X$ that
$u(y)=\langle u,y\rangle=\langle x,y\rangle_{X}.$ (7)
Note carefully that the notation $\langle u,y\rangle$, which we will also employ in the Banach-space case from here on, describes the dual pairing of $u\in X^{*}$ with $y\in X$ (i.e., the application of the linear functional $u$ to $y$), and should not be confused with $\langle x,y\rangle_{X}$, which is the inner product of elements in the Hilbert space $X$. This notation fits well with the previously mentioned case $X=L^{p}(\Omega)$ on some domain $\Omega\subseteq\mathbb{R}^{n}$, $1<p<\infty$, where $X^{*}=L^{p^{*}}(\Omega)$ with $1/p+1/p^{*}=1$, and
$u(x)=\langle u,x\rangle=\int_{\Omega}u(\bm{r})x(\bm{r})\mathrm{d}\bm{r},$ (8)
which equals $\langle u,x\rangle_{L^{2}}$ for $p=2$. So, in this case, we arrive back at the inner product on the Hilbert space $L^{2}(\Omega)$, which is called a pivot space between $L^{p}(\Omega)$ and $L^{p^{*}}(\Omega)$. In other words, the inner product of the pivot space serves as the duality pairing.
This construction can be generalized to arbitrary $X$ and $X^{*}$ within the formalism of a rigged Hilbert space (Gel’fand triple). Applied to the Hilbert space of wavefunctions, this approach provides a rigorous foundation for Dirac’s bra–ket formalism and can be employed to describe continuous spectra and scattering states in quantum mechanics [21; 22]. Those are different objects than the densities and potentials considered here, but we find the similarities still striking.
Note that the three assumption sets are increasingly strong—that is, Assumption 3 implies Assumption 2, which in turn implies our minimal requirement Assumption 1.
II.2 Convex analysis preliminaries
The second mathematical pillar of DFT is without any doubt convex analysis, since DFT admits a beautiful formulation in terms of convex and concave functionals on Banach spaces. On a Banach space $X$ that fulfills Assumption 1, we now consider convex functions $f\colon X\to\mathbb{R}\cup\{+\infty\}$ that are not $\infty$ everywhere. We also usually assume lower semicontinuity, which means that for each $x_{0}\in X$ it holds that $\liminf_{x\to x_{0}}f(x)\geq f(x_{0})$. Since the argument $x\in X$ of $f(x)$ is the element of a Banach space of functions that gets mapped to the scalar field, $f$ is sometimes called a functional (especially in the context of DFT). The Legendre–Fenchel transform (convex conjugate) of any $f:X\to\mathbb{R}\cup\{-\infty,+\infty\}$ is a function $f^{*}\colon X^{*}\to\mathbb{R}\cup\{-\infty,+\infty\}$ given by
$f^{*}(u)=\sup_{x\in X}\{\langle u,x\rangle-f(x)\}.$ (9)
The convex conjugate of a convex and lower semicontinuous $f\colon X\to\mathbb{R}\cup\{+\infty\}$ is again convex and lower semicontinuous. By applying this transformation twice and restricting the result again to $X\subseteq X^{**}$ (biconjugate)
$f^{**}(x)=\sup_{u\in X^{*}}\{\langle u,x\rangle-f^{*}(u)\}$ (10)
one arrives back at the original function, $f^{**}=f$, if and only if $f$ is convex and lower semicontinuous.
The subdifferential of $f$ at some $x\in X$ is the set-valued map
$\partial f(x)=\{u\in X^{*}\mid\forall y\in X\colon f(y)\geq f(x)+\langle u,y-x\rangle\}.$ (11)
On the real axis, the set $\partial f(x)$ can be interpreted as the collection of all slopes for which the corresponding tangent of $f$ at $x$ never exceeds the function $f$ itself (globally, since the function is assumed convex). It can be defined for concave functions as the superdifferential by just reverting the inequality, and we use the same notation in this case. Note that $\partial f(x)$ is always a convex set, and that it includes infinitely many elements where $f$ has a kink and is thus non-differentiable. It can also be the empty set. To illustrate this, take $X=\mathbb{R}$ and $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=+\infty$ for $x<0$ (which actually connects to reduced-density-matrix functional theory [23]) and then note that $\partial f(0)=\emptyset$. Another one-dimensional example can be taken from a DFT-treatment of the quantum Rabi model [24], where $\partial f(\pm 1)=\emptyset$ for $f(x)=-\sqrt{1-x^{2}}$ for $x\in[-1,1]$ and $f(x)=+\infty$ outside. If, on the other hand, $f$ is differentiable at $x$, then the subdifferential $\partial f(x)$ is single-valued and consists only of the functional derivative $\nabla f(x)$. The subdifferential fulfills the usual role of a derivative, by qualifying the minimum of a convex (not necessarily lower semicontinuous) functional,
$0\in\partial f(x_{*})\iff x_{*}\in\mathop{\mathrm{arg\,min}}_{x\in X}f(x).$ (12)
Yet, unlike the derivative, the subdifferential does not require the function to be differentiable. For a function $g$ on the dual space $X^{\ast}$, we define the subdifferential at $u\in X^{\ast}$ as a convex subset of the predual $X$ rather than the bidual $X^{\ast\ast}$,
$\partial g(u)=\{x\in X\mid\forall v\in X^{*}\colon g(v)\geq g(u)+\langle v-u,x\rangle\}.$ (13)
With these definitions in place, and if $f$ is lower semicontinuous, the subdifferentials of a pair of conjugate functions $(f,f^{*})$ are connected in the following way (inverse subdifferential relation),
$(\partial f)^{-1}=\partial f^{*},$ (14)
where the inverse is meant as a set-valued mapping.
II.3 Moreau–Yosida regularization
In order to deal with problems arising from a possible non-differentiability of $f\colon X\to\mathbb{R}\cup\{+\infty\}$ , we introduce the MY regularization with smoothing parameter $\varepsilon>0$ of a convex and lower semicontinuous functional $f$ as
$f_{\varepsilon}(x)=\inf_{y\in X}\left\{f(y)+\frac{1}{2\varepsilon}\|x-y\|_{X}^{2}\right\}.$ (15)
How such a MY regularized functional looks like is shown with a one-dimensional example of a DFT functional in Figure 1 in Section III.1. This construction has a long list of favorable properties that we reproduces here, while the proofs can be found in standard textbooks [5; 7]. It is important to remember that we always demand at least Assumption 1, ensuring that $X$ and $X^{\ast}$ are both reflexive and strictly convex:
  1. (i)
    $f_{\varepsilon}$ is convex, finite, and Gâteaux differentiable. If in addition $X^{*}$ is uniformly convex (Assumption 2) then $f_{\varepsilon}$ is also Fréchet differentiable.
  2. (ii)
    For each $x\in X$ and all $\varepsilon,\varepsilon^{\prime}$ such that $0<\varepsilon^{\prime}<\varepsilon$, we have $\inf f(X)\leq f_{\varepsilon}(x)\leq f_{\varepsilon^{\prime}}(x)\leq f(x)$ . In particular, $\inf f(X)=\inf f_{\varepsilon}(X)=\inf f_{\varepsilon^{\prime}}(X)$. The minimal value and the minimizers, if they exist, are preserved by the regularization.
  3. (iii)
    $\lim_{\varepsilon\to 0}f_{\varepsilon}(x)=f(x)$ for every $x\in X$.
  4. (iv)
    The infimum in Eq. 15 is attained at a unique point (proximal point) that defines the proximal map
    $\qquad\mathop{\mathrm{prox}}_{\varepsilon f}:X\to X,$ (16)
    $\qquad\mathop{\mathrm{prox}}_{\varepsilon f}(x)=\mathop{\mathrm{arg\,min}}_{y\in X}\left\{f(y)+\frac{1}{2\varepsilon}\|x-y\|_{X}^{2}\right\},$
    and $\lim_{\varepsilon\to 0}\mathop{\mathrm{prox}}_{\varepsilon f}(x)=x$ if $f(x)<\infty$.
  5. (v)
    $\nabla f_{\varepsilon}(x)=\varepsilon^{-1}J(x-\mathop{\mathrm{prox}}_{\varepsilon f}(x))\in\partial f(\mathop{\mathrm{prox}}_{\varepsilon f}(x))$.
  6. (vi)
    $f_{\varepsilon}(x)=f(\mathop{\mathrm{prox}}_{\varepsilon f}(x))+\frac{\varepsilon}{2}\|\nabla f_{\varepsilon}(x)\|_{X^{*}}^{2}$.
  7. (vii)
    For each $x\in X$ with $\partial f(x)$ non-empty, it holds that $\lim_{\varepsilon\to 0}\nabla f_{\varepsilon}(x)=w$ (weakly) where $w\in X^{*}$ is the (unique) element in $\partial f(x)$ with minimal norm. If $X^{*}$ is uniformly convex (Assumption 2), then the convergence is strong.
  8. (viii)
    For the Legendre–Fenchel transform of $f_{\varepsilon}$, it holds that
    $(f_{\varepsilon})^{*}(u)=f^{*}(u)+\tfrac{\varepsilon}{2}\|u\|_{X^{*}}^{2}\quad\text{and}$ (17)
    $\partial(f_{\varepsilon})^{*}(u)=\partial f^{*}(u)+\varepsilon J^{-1}(u)$ (18)
    for each $u\in X^{*}$. By adding $\tfrac{\varepsilon}{2}\|u\|_{X^{*}}^{2}$ to $f^{*}(u)$ the $(f_{\varepsilon})^{*}$ automatically becomes a strictly convex function.
Note that $X$ strictly convex was already needed for the definition of the proximal map since it means that $x\mapsto\|x\|^{2}_{X}$ is strictly convex and thus the minimum in Eq. (16) is reached at a unique point. The reason that we can write $\mathop{\mathrm{prox}}_{\varepsilon f}$ instead of $\mathop{\mathrm{prox}}_{f}^{\varepsilon}$ (or similar) in order to include the regularization parameter $\varepsilon$ in the notation is that the definition of the proximal map could have also been written as
$\mathop{\mathrm{prox}}_{\varepsilon f}(x)=\mathop{\mathrm{arg\,min}}_{y\in X}\left\{\varepsilon f(y)+\frac{1}{2}\|x-y\|_{X}^{2}\right\}.$ (19)
In a Hilbert-space setting (Assumption 3), a few additional useful properties of MY regularization and the proximal map can be added [6]:
  1. (ix)
    $f_{\varepsilon}$ is Fréchet differentiable and its functional derivative $\nabla f_{\varepsilon}$ is Lipschitz-continuous with constant $\varepsilon^{-1}$.
  2. (x)
    The proximal map is given by the resolvent of the subdifferential, $\mathop{\mathrm{prox}}_{\varepsilon f}(x)=(\mathop{\mathrm{id}}+\varepsilon\partial f)^{-1}(x)$.
  3. (xi)
    The proximal map is firmly nonexpansive, i.e., for all $x,y\in X$ it holds
    $\langle\mathop{\mathrm{prox}}_{\varepsilon f}(x)-\mathop{\mathrm{prox}}_{\varepsilon f}(y),x-y\rangle_{X}$ (20)
    $\qquad\geq\|\mathop{\mathrm{prox}}_{\varepsilon f}(x)-\mathop{\mathrm{prox}}_{\varepsilon f}(y)\|^{2}_{X}.$
    The same property holds for $\mathop{\mathrm{id}}-\mathop{\mathrm{prox}}_{\varepsilon f}$.
  4. (xii)
    $\partial(f_{\varepsilon})^{*}$ is strongly monotonous, i.e., for all $u,v\in X^{*}$ and $x\in\partial(f_{\varepsilon})^{*}(u)$, $y\in\partial(f_{\varepsilon})^{*}(v)$, it holds
    $\langle u-v,x-y\rangle\geq\varepsilon\|u-v\|^{2}_{X^{*}}.$ (21)
While the notion of firm nonexpansiveness itself can be generalized to Banach spaces [25], it does not necessarily hold for the proximal map anymore. Nevertheless, a result of uniform continuity can be achieved for the proximal map also in a certain Banach-space setting [26]. Combining v from the previous properties and the resolvent form for the proximal map yields
$\nabla f_{\varepsilon}(x)=\varepsilon^{-1}\left(x-(\mathop{\mathrm{id}}+\varepsilon\partial f)^{-1}(x)\right),$ (22)
which is called the Yosida approximation [6, Def. 23.1] of the multivalued mapping $\partial f$ in functional analysis. It was first published by Yosida around the same time as Moreau’s contributions to convex analysis and together they became the namesakes of the regularization method.
II.4 Formal density-functional theory
The standard setting of DFT is the $N$-particle Hamiltonian with an external, scalar potential $v$ on $\mathbb{R}^{3}$, where we use atomic units,
$\hat{H}^{\lambda}_{v}=\hat{T}+\lambda\hat{W}+\sum_{j=1}^{N}v(\bm{r}_{j}),$ (23)
$\hat{T}=-\frac{1}{2}\sum_{j=1}^{N}\nabla_{j}^{2},\quad\hat{W}=\sum_{j<k}|\bm{r}_{j}-\bm{r}_{k}|^{-1}.$
Here, we allow different choices of the interaction strength $\lambda$, which can range from the non-interacting case ($\lambda=0$) to the fully interacting case with a repulsive force ($\lambda=1$). For all choices of $\lambda\in\mathbb{R}$ and $v\in L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$, this Hamiltonian is self-adjoint and bounded below on its usual domain [27, Th. X.19]. It was already mentioned that the methods and results presented here readily apply to various settings outside of standard DFT, although subtleties around the involved spaces and definitions of the basic functionals have to be taken into account. A setting that also involves the vector potential is briefly discussed in Section VI.2.
The ground-state problem of quantum mechanics is given by the Rayleigh–Ritz variational principle,
[…]
where the variation is either over all normalized wave functions or density matrices with finite kinetic energy and the required antisymmetry. The result, $E^{\lambda}(v)$, is then the ground-state energy of the system for a given external potential $v$ and chosen interaction strength. The principal idea of DFT is now to separate a density-dependent part inside this variation, which is universal in the sense that it does not depend on $v$ [28; 29; 30; 3]. An elegant way to do this is by Levy’s constrained search—that is, one varies over all wave functions or density matrices that yield a given one-particle density
$\rho_{\Psi}(\bm{r})$ $=N\int|\Psi(\bm{r},\bm{r}_{2},\ldots,\bm{r}_{N})|^{2}\mathrm{d}\bm{r}_{2}\ldots\mathrm{d}\bm{r}_{N},$ (25)
$\rho_{\hat{\Gamma}}(\bm{r})$ $=\sum_{j}p_{j}\rho_{\Psi_{j}}(\bm{r}),\quad p_{j}\in[0,1],\quad\sum_{j}p_{j}=1.$ (26)
We write $\Psi\mapsto\rho$ for an $N$-particle wave function that has $\rho_{\Psi}=\rho$ and analogously for density matrices. We define the universal density functional by a constrained search over pure states and ensembles, respectively,
$\tilde{F}^{\lambda}(\rho)=\inf_{\Psi\mapsto\rho}\langle\Psi,\hat{H}^{\lambda}_{0}\Psi\rangle,$ (27)
[…]
where $\hat{H}^{\lambda}_{0}=\hat{T}+\lambda\hat{W}$. Since the constrained search over ensembles is over a larger set, we have $F_{\mathrm{DM}}^{\lambda}(\rho)\leq\tilde{F}^{\lambda}(\rho)$. The functional $F_{\mathrm{DM}}^{\lambda}(\rho)$ is convex since it is given as the infimum of a linear functional over a convex set, unlike $\tilde{F}^{\lambda}(\rho)$, which is therefore not necessarily convex whenever $\tilde{F}^{\lambda}(\rho)\neq F_{\mathrm{DM}}^{\lambda}(\rho)$ [31; 3], and explicit non-convex examples are known [32, Sec. VI.E].
The effective domain of a (convex) functional is the set of points where it is finite. For the constrained-search functionals $\tilde{F}^{\lambda}(\rho)$ and $F_{\mathrm{DM}}^{\lambda}(\rho)$, the effective domains agree and are also independent of $\lambda$ and we set
$D=\{x\in X\mid\tilde{F}^{\lambda}(x)<\infty\}=\{x\in X\mid F_{\mathrm{DM}}^{\lambda}(x)<\infty\},$ (29)
which is closed if the convex $F_{\mathrm{DM}}^{\lambda}(x)$ is lower semicontinuous. From the definition of the density $\rho_{\Psi}$ and the definition of the functionals it follows that any $\rho\in D$ must be non-negative and normalized to the number of particles. It is a fact that $D$ corresponds to the $N$-representable densities [3, Th. 1.2]. By definition we have $\tilde{F}^{\lambda}(x)=F_{\mathrm{DM}}^{\lambda}(x)=+\infty$ for $x\in X\setminus D$. For this reason, the usual convention will be to use the symbol $\rho$ for elements of $D$ that are ‘real’ densities, and use $x,y\in X$ otherwise.
The pure- and ensemble-state density functionals given in Eqs. (27) and (28), respectively, yield the same ground-state energy by the Hohenberg–Kohn variation principle,
$E^{\lambda}(v)$ $=\inf_{\rho\in X}\{\tilde{F}^{\lambda}(\rho)+\langle v,\rho\rangle\}$ (30)
$=\inf_{\rho\in X}\{F_{\mathrm{DM}}^{\lambda}(\rho)+\langle v,\rho\rangle\},$
where the dual pairing of potential and density yields the potential energy, $\langle v,\rho\rangle=\int v(\bm{r})\rho(\bm{r})\mathrm{d}\bm{r}$. Recognizing that Eq. (30) has the structure of the Legendre–Fenchel transform introduced in Section II.2, we define the back transform
$F^{\lambda}(\rho)=(-E^{\lambda})^{*}(-\rho)\leq F_{\mathrm{DM}}^{\lambda}(\rho)$ (31)
as a new functional that is convex and lower semicontinuous by construction. We can then get $E^{\lambda}(v)$ from this functional as well,
$E^{\lambda}(v)=-(F^{\lambda})^{*}(-v).$ (32)
Now if the convex $F_{\mathrm{DM}}^{\lambda}(\rho)$ is also lower semicontinuous (Assumption 4 below), this implies
$F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho).$ (33)
Unlike convexity, the lower semicontinuity of $F_{\mathrm{DM}}^{\lambda}(\rho)$ does not follow automatically and needs to be checked for every choice of topology. In settings where $F_{\mathrm{DM}}^{\lambda}(\rho)$ is not lower-semicontinuous, the convex conjugate of $E^{\lambda}(v)$ does notably not retrieve the original constrained-search functional $F_{\mathrm{DM}}^{\lambda}(\rho)$ in Eq. (31). Thus it looses connection to the variational principle in Eq. (24) and consequently also to the equivalent ground-state problem involving the Schrödinger equation. This means that minimizers $\rho\in X$ in Eq. (30) can occur where the convex conjugate $F^{\lambda}(\rho)$ of $E^{\lambda}(v)$ is finite, yet the constrained-search functional $F_{\mathrm{DM}}^{\lambda}(\rho)=+\infty$. If, on the other hand, $v$ in $E^{\lambda}(v)$ supports a ground state $\Psi$, then $F_{\mathrm{DM}}^{\lambda}(\rho_{\Psi})$ is also finite. For this reason it needs to be stressed that albeit the requirement of lower semicontinuity for the ensemble constrained-search functional seems like a mathematical sophistry, it is a critical feature of the theory, and we will elaborate further on this on page 39. The convex $F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho)$ on $X$ and the concave $E^{\lambda}(v)$ on $X^{*}$, as a conjugate pair, then serve as the backbone of a convex-analysis formulation of DFT.
Based on the above discussion, we formulate a separate assumption set for a consistent DFT framework.
Assumption 4.
$X$ and $X^{*}$ must be such that the Hamiltonian $\hat{H}_{v}^{\lambda}$ is self-adjoint with a $v$-independent domain and bounded below, and the ensemble constrained-search functional $F_{\mathrm{DM}}^{\lambda}(\rho)$ is lower semicontinuous. This implies that the $N$-representability set $D$ is closed and $F_{\mathrm{DM}}^{\lambda}(\rho)=F^{\lambda}(\rho)$.
In order to fulfill this assumption, we can think about Lieb’s choice $L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$ for the space of densities $X$, which indeed makes $F_{\mathrm{DM}}^{\lambda}(\rho)$ lower semicontinuous [3]. Yet, this space does not fulfill even Assumption 1, which is also not necessary for a consistent formulation of DFT. However, this assumption is needed for our later constructions involving MY regularization. Instead, we suggest to switch to a more general yet reflexive $X\supsetneq L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$. This, on the other hand, can break lower semicontinuity of the ensemble constrained-search functional if we switch to a coarser topology.
The dual pairing between potentials and densities in $X$ suggests a matching space for potentials that consists of the topological dual $X^{*}$. For $L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$ the dual space is $L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$. If instead $X\supsetneq L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$, then $X^{*}\subsetneq L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$. We have to be careful that this smaller dual space still allows for a self-adjoint and bounded-below Hamiltonian on its usual domain, but indeed a result for such potentials is available [27, Th. X.19]. Interestingly, $L^{3/2}(\mathbb{R}^{3})$ is exactly the boundary case in terms of $L^{p}$ spaces and the Hamiltonian fails to be self-adjoint on its usual domain for $p<3/2$ when the potential can become too singular. This is the case for a potential $r^{-\alpha}$, $\alpha\geq 2$, where different self-adjoint realizations on the punctured domain $\mathbb{R}^{3}\setminus\{0\}$ become possible. To ensure a self-adjoint and bounded-below Hamiltonian on the usual domain, we thus need to limit the potential space to $X^{*}\subset L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$.
At this point, another problem appears due to this limitation. In certain cases, this space will not accommodate a given physical setting—for example, it might exclude singular Coulomb potentials and thus molecular systems in a clamped nuclei or Born–Oppenheimer approximation. We can solve this problem by just including a fixed offset potential from a more general class,
$v_{\mathrm{o}}\in L^{2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})\subsetneq L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3}).$ (34)
It can also be taken from any other potential class that still admits a self-adjoint and bounded-below Hamiltonian according to Assumption 4, including non-local potentials. In particular, $v_{\mathrm{o}}$ can fulfill the task of a binding potential for $N$ electrons. But one has to take careful note that such a potential can change the domain of the self-adjoint Hamiltonian if it is not from the class of Eq. (34) [27, Sec. X.2]. Instead of writing $E^{\lambda}(v_{\mathrm{o}}+v)$, we redefine the energy functional as
$E^{\lambda}(v)=\inf_{\Psi}\langle\Psi,\hat{H}^{\lambda}_{v_{\mathrm{o}}+v}\Psi\rangle,$ (35)
including the fixed offset automatically. Note that this also affects the universal density functional as the corresponding Legendre–Fenchel transform that can be retrieved as
[…]
While the inclusion of this offset potential does not change anything within the basic structure of DFT, which is the reason for not including it explicitly into our notation, it cannot be ignored when solving the underlying equations, and we will highlight its occurrence at several places in the following sections.
Whenever a minimizer can be found in the Hohenberg–Kohn variation principle of Eq. (30) with the functional $F_{\mathrm{DM}}^{\lambda}(\rho)$, we call this $\rho\in D$ a $v$-representable density. This means, in other words, that there is a $v\in X^{*}$ such that the (ensemble) ground-state solution $\hat{\Gamma}$ to the Schrödinger equation with Hamiltonian $\hat{H}^{\lambda}_{v_{\mathrm{o}}+v}$ yields $\hat{\Gamma}\mapsto\rho$ and ground-state energy $E^{\lambda}(v)=F_{\mathrm{DM}}^{\lambda}(\rho)+\langle v,\rho\rangle$. We next write out the relation in Eq. (31) as the Lieb variational principle,
$F^{\lambda}(\rho)=\sup_{v\in X^{*}}\{E^{\lambda}(v)-\langle v,\rho\rangle\}.$ (37)
Since in general $F^{\lambda}(\rho)\leq F_{\mathrm{DM}}^{\lambda}(\rho)$ and since a $v$-representable $\rho$ is a minimizer in Eq. (30) and must fulfill $F_{\mathrm{DM}}^{\lambda}(\rho)=E^{\lambda}(v)-\langle v,\rho\rangle$, we see that $F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho)$ always holds for $v$-representable densities, even when $F_{\mathrm{DM}}^{\lambda}(\rho)$ is not lower semicontinuous. In what follows we want to relate this property to the functional $F^{\lambda}(\rho)$ and its subdifferential.
The notions of sub- and superdifferentials serve an important purpose in the formulation of DFT. If there exists an optimizer $v\in X^{*}$ for a given $\rho\in X$ in Eq. (37), then the following are all equivalent for the lower-semicontinuous $F^{\lambda}(\rho)$,
$v\in-\partial F^{\lambda}(\rho)$ $\iff\rho\in\partial E^{\lambda}(v)$ (38)
$\iff E^{\lambda}(v)=F^{\lambda}(\rho)+\langle v,\rho\rangle.$
We can combine this into the inverse subdifferential relation,
$(-\partial F^{\lambda})^{-1}(v)=\partial E^{\lambda}(v).$ (39)
Yet, it is by no means secured that the set $\partial F^{\lambda}(\rho)$ is non-empty for all $\rho\in D$. But since $F^{\lambda}(\rho)$ is convex and lower semicontinuous by definition, it is known that its subdifferential is non-empty on a dense subset of the effective domain of the functional [5, Cor. 2.44]. But only if $F_{\mathrm{DM}}^{\lambda}(\rho)$ is lower semicontinuous too, we have $F_{\mathrm{DM}}^{\lambda}(\rho)=F^{\lambda}(\rho)$, and what was said above also applies to $F_{\mathrm{DM}}^{\lambda}(\rho)$. In addition, in this case, any $\rho\in\partial E^{\lambda}(v)$ is an optimizer in the second line of Eq. (30) and therefore also $\rho\in D$ . The convex set $\partial E^{\lambda}(v)\subseteq D$ then yields all possible ground-state densities for the potential $v\in X^{*}$. This $\partial E^{\lambda}(v)$ will be empty if $v$ does not support a ground state for interaction strength $\lambda$, includes a single element if $v$ supports a nondegenerate ground state, and contains multiple elements in case of degeneracy. We may now identify the points at which $F^{\lambda}(\rho)$ has a non-empty subdifferential with the ($v$-representable) ground-state densities of a Schrödinger equation. Hence, assuming a lower semi-continuous $F_{\mathrm{DM}}^{\lambda}(\rho)$, it also holds that the set of $v$-representable densities is dense in the set of $N$-representable densities. This way, it can be highlighted in a different way why the lower semicontinuity of the ensemble constrained-search functional is important. Else, with a $F_{\mathrm{DM}}^{\lambda}(\rho)$ that is not lower semicontinuous, it is possible to find $v\in-\partial F^{\lambda}(\rho)$ when at the same time $F_{\mathrm{DM}}^{\lambda}(\rho)=+\infty$, so this density has $\rho\in X\setminus D$ and would be clearly not $v$-representable. This is why Assumption 4 is important when dealing with the functionals $F_{\mathrm{DM}}^{\lambda}(\rho)$ $F^{\lambda}(\rho)$. See also the example given on page II.4.
But it is also known since a long time that $v$-representability, even if it is secured for a dense set, is a problematic notion since many densities $\rho\in D$ will not have this property [33; 34]. Even worse, in the standard setting of DFT with density space $L^{1}\cap L^{3}$, arbitrarily close to any $v$-representable density there lies a non-$v$-representable one, which makes $F^{\lambda}(\rho)$ discontinuous and non-differentiable [35]. In many cases (including standard DFT) the set of $v$-representable densities is known only implicitly (exceptions are lattice systems [36; 32; 37] and particles on a one-dimensional torus [17; 18]). We will later see how MY regularization offers a way out of this dilemma.
The Hohenberg–Kohn theorem [28] states that, for a $v$-representable density $\rho\in D$, the corresponding potential is unique up to addition of constants. In standard DFT, a rigorous proof of the Hohenberg–Kohn theorem has been achieved for $v\in L^{p}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$, with $p>2$ [38] and it extends to closely related settings [39]. This leaves $3/2\leq p\leq 2$ as undecided (the lower bound is given by the argument for self-adjoint Hamiltonians as noted before), although results on unique continuation [40; 41] suggest that the Hohenberg–Kohn theorem holds also in this range. In non-standard DFT settings, like finite lattices and when currents are included, counterexamples are known [32; 42].
Combining the assumptions from this section and from Section II.1 (which will later serve as the basis for MY regularized DFT), we arrive at the following important result:
Proposition. Assumption 1 together with Assumption 4 with a potential space $X^{*}\subset L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ imply that every $v\in X^{*}$ supports a ground state—that is, for all $v\in X^{*}$, the superdifferential $\partial E^{\lambda}(v)$ is non-empty.
This claim can be seen as follows. The choice $L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ as a potential space makes $E^{\lambda}(v)$ continuous [3, Th. 3.1(iii)], implying that each continuously embedded $X^{*}\subset L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ gives a continuous energy functional $E^{\lambda}(v)$. But a continuous and concave functional on $X^{*}$ admits a non-empty subdifferential belonging to $X^{\ast\ast}$ everywhere in the interior of its domain (here all $X^{\ast}$) [5, Prop. 2.36]. In general, it only holds $X^{**}\supseteq X$, but with Assumption 1 we have reflexivity and thus $X=X^{**}$. By the additional requirement that $F_{\mathrm{DM}}^{\lambda}(\rho)$ is lower semicontinuous, we have $F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho)$. Each $\rho\in\partial E^{\lambda}(v)$ is now in $D\subset X$ and through $v\in-\partial F^{\lambda}(\rho)=-\partial F_{\mathrm{DM}}^{\lambda}(\rho)$ it is connected to an actual ground-state solution of the Schrödinger equation as argued above. This demonstrates the claim of the proposition.
We illustrate this subtle and important point with an example. Take the Hamiltonian from Eq. (23) with just a single particle and set $v=0$. Then there is no bound state and thus $\partial E^{\lambda}(v)$ should be empty (note that the $\lambda$ parameter is not needed for single particles since there is no interaction anyway). Indeed, this example is incompatible with reflexivity and lower semicontinuity assumed together. If we give up on reflexivity of the spaces, then we can find an $x\in\partial E^{\lambda}(0)$ where $x\in X^{**}\setminus X$, so it does not represent a ground state. Next, we keep reflexivity but give up on lower semicontinuity of $F_{\mathrm{DM}}(\rho)$. We pick a minimizing sequence $\rho_{i}\to x$ in Eq. (30) consisting, for example, normalized Gaussian-shaped densities that get broader with increasing $i$ and are all in $D$, while $x\neq D$. Note that this means $x=0$, which cannot hold if $X$ includes the $L^{1}$ norm (and which would make it non-reflexive), since all $\rho_{i}$ were assumed normalized to the same number and clearly $N=\|\rho_{i}\|_{L^{1}}\to 0$ is a contradiction. But it is not necessarily a contradiction in a reflexive $X$ without the $L^{1}$ norm. Now the chosen sequence fulfills $\lim F_{\mathrm{DM}}(\rho_{i})=0$ but at the same time $F_{\mathrm{DM}}(x)=+\infty$, since $x\notin D$, so lower semicontinuity cannot hold.
This section summarized the most important features of DFT in the language of functional and convex analysis. The reader may wonder if Assumption 1 (or one of the two stronger ones) and Assumption 4 can be satisfied simultaneously. After all, the classical setting of DFT, with $X=L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$ for densities and its dual $X^{\ast}=L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ for potentials, allows for Assumption 4 but breaks reflexivity, a property not needed for a consistent formulation of DFT but necessary for our later results related to Moreau–Yosida regularization. One solution is to choose a bounded domain $\Omega$ which implies a continuous embedding $L^{p}(\Omega)\subseteq L^{q}(\Omega)$, $q<p$. Given that the important result of lower semicontinuity due to Lieb [3, Th. 4.4 and Cor. 4.5] holds with respect to the $L^{1}$ topology, it then extends to each $L^{p}(\Omega)$, $p>1$. We can then choose any $X=L^{p}(\Omega)$ with $1<p<\infty$ with dual $X^{*}=L^{p^{*}}(\Omega)$, $1/p+1/p^{*}=1$, and thus have Assumptions 2 and 4 simultaneously satisfied. By selecting $p=2$, we satisfy both Assumptions 3 and 4. The previous works of Eschrig [13] and Kvaal et al. [1] used this route, while many other works on the topic had to give up on at least one of the assumptions. Rather than switching to a bounded domain, we can use the offset potential $v_{\mathrm{o}}$ to introduce a confining potential for the whole system and thus ensure that every $v\in X^{*}$ yields a ground-state density. This, at the same time, serves to break the full translational and rotational symmetry of the setting, which is necessary to allow for any localized ground state in the first place [43].
II.5 Kohn–Sham method
DFT reformulates the ground-state problem of quantum mechanics by replacing the many-particle wave function as the central variable with the one-particle density, thereby significantly reducing the complexity of the description. Yet, the usefulness of such a reduced description entirely depends on whether $F^{\lambda}(\rho)$ can be evaluated accurately and efficiently. The $F^{0}(\rho)$ of a non-interacting system amounts to just the kinetic energy and its conjugate functional $E^{0}(v)$ can be computed effectively by means of orbitals. For the interacting case, however, $F^{1}(\rho)$ is in principle unavailable since its evaluation corresponds to a quantum version of an NP-hard problem [44]. The difference $F^{1}(\rho)-F^{0}(\rho)$, which relates to the exchange–correlation functional, has moreover turned out to be difficult to model in terms of $\rho$ [45]. Thus, although the exact DFT formulation is very appealing by avoiding the full description by a wave function, further steps are needed to attain a useful reformulation. In the KS method [46], an auxiliary system of non-interacting electrons is introduced and its external potential $v_{\mathrm{s}}$ is chosen in such a way that it reproduces the same density as the interacting system. Yet, we will argue below that it is by no means clear that this approach is always possible.
The exchange–correlation potential is defined as the potential that must be added to the external potential and the Hartree potential of the auxiliary, non-interacting system in order to get the same ground-state density as the interacting system. This means that the method requires $\rho\in D$ to be simultaneously interacting and non-interacting $v$-representable, which is equivalent to $\partial F^{1}(\rho)$ and $\partial F^{0}(\rho)$ both non-empty. We already noted on the general issues regarding $v$-representability, and this additional requirement of simultaneous $v$-representability comes as a further restriction [47]. Assuming for now that this is ensured, we can take $v\in-\partial F^{1}(\rho)$ and $v_{\mathrm{s}}\in-\partial F^{0}(\rho)$ and define the exchange–correlation potential
$v_{\mathrm{xc}}(\rho)=v_{\mathrm{s}}-v-v_{\mathrm{H}}(\rho).$ (40)
Here we already subtracted the Hartree potential
$v_{\mathrm{H}}(\rho)=\nabla E_{\rm H}(\rho)=\int\frac{\rho(\bm{r}^{\prime})}{|\cdot-\bm{r}^{\prime}|}\mathrm{d}\bm{r}^{\prime},$ (41)
that itself is the functional derivative of the Hartree mean-field energy
$E_{\rm H}(\rho)=\frac{1}{2}\iint\frac{\rho(\bm{r})\rho(\bm{r}^{\prime})}{|\bm{r}-\bm{r}^{\prime}|}\mathrm{d}\bm{r}\mathrm{d}\bm{r}^{\prime}.$ (42)
The total external potential that needs to be applied to the non-interacting system in order to get $\rho$ is then, not forgetting the previously introduced offset potential,
$v_{\mathrm{o}}+v_{\mathrm{s}}=v_{\mathrm{o}}+v+v_{\mathrm{H}}(\rho)+v_{\mathrm{xc}}(\rho),$ (43)
As was already noted, the assumption of (simultaneous interacting and non-interacting) $v$-representability comes as a severe limitation. But putting this issue aside for the moment and assuming the existence of $v_{\mathrm{xc}}(\rho)$ for all $\rho$ in question, Eq. (43) can be translated into the KS self-consistent iteration scheme,
$\begin{array}[]{cccccc}v_{\mathrm{s}}&=v+v_{\mathrm{H}}(&\rho&)+v_{\mathrm{xc}}(&\rho&)\\ \updownarrow&&\updownarrow&&\updownarrow&\\ v_{i+1}&=v+v_{\mathrm{H}}(&\rho_{i}&)+v_{\mathrm{xc}}(&\rho_{i}&)\end{array}$ (44)
and
$\rho_{i+1}\in\partial E^{0}(v_{i+1}).$ (45)
The second step corresponds to solving the non-interacting Schrödinger equation with an external potential $v_{i+1}$ and determining a ground-state density $\rho_{i+1}$. Since the Hamiltonian is bounded below for $v_{i+1}\in X^{*}$, this approach is always possible. A more detailed convex-analytical argument for the non-emptiness of the superdifferential can be found elsewhere [14, proof of Th. 12]. The two steps of Eqs. (44) and (45) can be combined into a single step, either for the potentials or, as we will show here, for the densities,
$\rho_{i+1}\in\partial E^{0}(v+v_{\mathrm{H}}(\rho_{i})+v_{\mathrm{xc}}(\rho_{i})).$ (46)
For initialization, we choose an initial density $\rho_{1}$ or, equivalently, a potential $v_{1}$ for the auxiliary system that roughly leads to the desired ground-state density $\rho$. Since the auxiliary system does not include any interactions between the particles, we (in the simplest case) choose its ground state as a Slater determinant and solve Eq. (45) by filling up the orbitals of a one-particle Schrödinger equation with external potential $v_{\mathrm{o}}+v_{i+1}$. If the procedure converges, we get
$\rho_{i}\to\rho,\qquad v_{i}\to v_{\mathrm{s}},$ (47)
so indeed both systems share the same ground-state density if in the auxiliary system an external potential $v_{\mathrm{o}}+v_{\mathrm{s}}$ is applied. Finally, in order to conveniently retrieve the interacting energy $E^{1}(v)$ from a converged density and the expressions defined above, we can just use $E^{0}(v_{\mathrm{s}})=F^{0}(\rho)+\langle v_{\mathrm{s}},\rho\rangle$ and calculate
$E^{1}(v)$ $=F^{1}(\rho)+\langle v,\rho\rangle$ (48)
$=F^{0}(\rho)+F^{1}(\rho)-F^{0}(\rho)+\langle v-v_{\mathrm{s}}+v_{\mathrm{s}},\rho\rangle$
$=E^{0}(v_{\mathrm{s}})+E_{\mathrm{xc}}(\rho)+E_{\rm H}(\rho)-\langle v_{\mathrm{xc}}(\rho)+v_{\mathrm{H}}(\rho),\rho\rangle$
$=E^{0}(v_{\mathrm{s}})+E_{\mathrm{xc}}(\rho)-E_{\rm H}(\rho)-\langle v_{\mathrm{xc}}(\rho),\rho\rangle,$
where we have introduced the exchange–correlation functional $E_{\mathrm{xc}}(\rho)=F^{1}(\rho)-F^{0}(\rho)-E_{\rm H}(\rho)$. We remark that practical KS-DFT methods are usually based on approximating $E_{\mathrm{xc}}(\rho)$. Assuming differentiability, we then have $v_{\mathrm{xc}}=\nabla E_{\mathrm{xc}}(\rho)$.
Note that this formulation of the KS method does not come with any guarantee of convergence. Indeed, in practice, an additional damping step [48; 49; 50] like taking $\rho_{i}+\alpha(\rho_{i+1}-\rho_{i})$, $\alpha<1$, as the next density is introduced to facilitate convergence. A mathematical proof of convergence has only been achieved in the MY-regularized setting, which is in any case needed to avoid the $v$-representability problem, as will be discussed in Section V.3.
III Moreau–Yosida regularization in density-functional theory
Having the stage set, we now show three different ways to introduce MY regularization into DFT, which are equivalent and thus have the same outcome—namely, replacing the universal density functional $F^{\lambda}(\rho)$ with a differentiable $F_{\varepsilon}^{\lambda}(x)$ for all $x\in X$. The following sections will achieve this task. In Section III.2, we show that $F_{\varepsilon}^{\lambda}(x)$ is the convex conjugate of an energy functional $E_{\varepsilon}^{\lambda}(v)$ that includes the field energy of $v$ (if appropriately defined), while Section III.3 identifies every $x\in X$ as a mix of internal and external densities. In the section we will only require Assumption 1.
III.1 MY based on the universal density functional
The most straightforward way to introduce MY regularization into DFT follows from its definition in Eq. (15): By directly applying the method to the non-differentiable (but convex and lower semicontinuous) $F^{\lambda}(\rho)$, yielding the regularized universal density functional $F_{\varepsilon}^{\lambda}(x)$, which is well-defined and finite for all $x\in X$,
$F_{\varepsilon}^{\lambda}(x)=\inf_{\rho\in X}\left\{F^{\lambda}(\rho)+\frac{1}{2\varepsilon}\|x-\rho\|_{X}^{2}\right\},$ (49)
where $\varepsilon>0$ is the smoothing parameter. An illustration of a regularized functional in one dimension together with the potentially non-convex $\tilde{F}^{\lambda}(x)$ and convex $F^{\lambda}(x)$ is shown in Figure 1. The regularized functional has a well-defined functional derivative (see v from the properties of MY regularization in Section II.3) and the regularization is lossless in the sense that it conserves the overall infimum of the functional (see ii). We next introduce the corresponding energy functional $E_{\varepsilon}^{\lambda}(v)$ as the convex conjugate of $F_{\varepsilon}^{\lambda}(x)$,
$E_{\varepsilon}^{\lambda}(v)=-(F_{\varepsilon}^{\lambda})^{*}(-v)=\inf_{x\in X}\{F_{\varepsilon}^{\lambda}(x)+\langle v,x\rangle\}.$ (50)
Note that the resulting $E_{\varepsilon}^{\lambda}(v)$ is not the MY regularization of $E^{\lambda}(v)$, but instead the following relation holds (from Eq. (17), where the different sign is due to Eq. (32)),
$E_{\varepsilon}^{\lambda}(v)=E^{\lambda}(v)-\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}.$ (51)
Since we subtract the strictly concave norm-square of $v$ from the concave $E^{\lambda}(v)$, the MY energy $E_{\varepsilon}^{\lambda}(v)$ is strictly concave. Conversely, it holds that the Legendre–Fenchel transform of a strictly concave energy functional is always differentiable. It is precisely this expression for the energy functional from where we start in the next motivation for MY regularization that is more physically informed.
With $v=-\nabla F_{\varepsilon}^{\lambda}(x)$ we now have a unique representing potential $v\in X^{*}$ for every $x\in X$. We have already written $\nabla F_{\varepsilon}^{\lambda}$ instead of the subdifferential $\partial F_{\varepsilon}^{\lambda}$ because we know that the regularized functional is differentiable. This then automatically yields a well-defined and unique mapping from ‘densities’ (elements of $X$) to potentials (elements of $X^{*}$) without any $v$-representability issue. We are not even limited to the $N$-representability set $D\subsetneq X$ any more, and Section III.3 will help to make sense of the generalized densities $x\in X$. Uniqueness of the potential also means we automatically have a Hohenberg–Kohn-type result.
In the Hilbert-space setting (Assumption 3), we obtain also a quantitative bound that shows how the difference between two potentials is bound by the difference between the corresponding densities, in form of the Lipschitz estimate from property ix,
$\|\nabla F_{\varepsilon}^{\lambda}(x)-\nabla F^{\lambda}_{\varepsilon}(y)\|_{X^{*}}\leq\varepsilon^{-1}\|x-y\|_{X}.$ (52)
This bound is obviously lost in the limit $\varepsilon\to 0$, in agreement with the fact that, in standard DFT, two densities that lie very close to each other can come from vastly different potentials [51], making density–potential inversion notoriously difficult. A similar bounds is also possible in more general Banach spaces if the MY regularization is adapted to the ‘modulus of convexity’ of the space [52]. Furthermore, MY regularization allows for a well-defined formulation of the KS iteration procedure (Section V).
$\tilde{F}^{\lambda}(x)$$F^{\lambda}(x)$$F_{\varepsilon}^{\lambda}(x)$
Figure 1: Example of a pure-state functional $\tilde{F}^{\lambda}(x)$, the convex $F^{\lambda}(x)$, and its MY regularization $F_{\varepsilon}^{\lambda}(x)$. The unregularized functionals jump to $\infty$ for all $x$ outside their effective domain $D$, but the regularized functional remains finite everywhere.
III.2 MY based on the total energy functional
A different viewpoint starts from the concave ground-state energy functional $E^{\lambda}(v)$ and introduces a strictly concave energy $E_{\varepsilon}^{\lambda}(v)$ by just subtracting $\varepsilon\phi(v)=\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$, i.e.,
$E_{\varepsilon}^{\lambda}(v)=E^{\lambda}(v)-\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}.$ (53)
The convex conjugate of this strictly concave functional then yields the regularized density functional of Eq. (49) as shown by Eq. (17),
$F_{\varepsilon}^{\lambda}(x)=(-E_{\varepsilon})^{*}(-x)=\sup_{v\in X^{*}}\left\{E_{\varepsilon}^{\lambda}(v)-\langle v,x\rangle\right\}.$ (54)
We have thus defined MY regularization starting from the energy functional instead of the universal density functional. But talking about the total energy allows to give this manipulation a physical interpretation. Say, for any external potential $v\in X^{*}$, we define the energy functional
$E^{\mathrm{f}}(v)=-\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}.$ (55)
Note the minus sign that makes this functional concave, just like the concave $E^{\lambda}(v)$ of the matter system. Of course, to retrieve the correct physical electrostatic field energy, $\frac{\varepsilon_{0}}{2}\|\nabla v\|^{2}_{L^{2}}$, an appropriate choice of Banach space $X^{*}$ must be made; see Section VI.1. We then introduce the new energy functional (that could be interpreted as the total energy of matter and field)
$E_{\varepsilon}^{\lambda}(v)=E^{\lambda}(v)+E^{\mathrm{f}}(v),$ (56)
which corresponds to Eq. (53). In this way, MY regularization arises naturally by attaching more physical relevance to the external potential and by assigning to it an energy that has to be taken into account in the total energy functional.
Regarding the negative sign in Eq. (55), remember that this is the (concave) energy functional and not the internal energy expression that is actually varied over. There, the additional term is $\tfrac{1}{2\varepsilon}\|x-\rho\|^{2}_{X}$ as seen in Eq. (49) and using the relation motivated in Eq. (57) below this is equal to $\tfrac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$, precisely the expression for the (positive) field energy. Viewed in this manner, it is the energy content of the external field that regularizes $F^{\lambda}(\rho)$. Conceptually, this is an extremely attractive viewpoint, since it explains the regularization as a physical effect.
III.3 MY based on density–potential mixing
Remember that we defined the $N$-representability set $D\subsetneq X$ in Eq. (29) as the set of all $\rho\in X$ where $F_{\mathrm{DM}}^{\lambda}(\rho)<\infty$, known as the effective domain of $F_{\mathrm{DM}}^{\lambda}(\rho)$. For the sake of argument, we now suppose Assumption 4 under which $F_{\mathrm{DM}}^{\lambda}(\rho)=F^{\lambda}(\rho)$. Now even for $\rho\in D$, we can have issues with $\partial F^{\lambda}(\rho)$. First, $\partial F^{\lambda}(\rho)$ may be empty. In such cases, no potential $v\in X^{\ast}$ exists such that $v\in-\partial F^{\lambda}(\rho)\iff\rho\in\partial E^{\lambda}(v)$, implying that $\rho$ is not $v$-representable. Also, a small variation of $\rho$ can (and will) lead outside of the set of $v$-representable densities [35; 51] and there are examples on small lattice systems where even large variations of $v$ lead to almost (or even exactly) the same ground-state density $\rho$ [32; 53].
It would be better altogether if we could ignore this difficult notion of “$v$-representability” and instead give a physical interpretation to all elements $x\in X$. This is achieved with the inverse duality map $J^{-1}\colon X^{*}\to X$ as the canonical way of mapping potentials to densities. For any $v$-representable $\rho\in\partial E^{\lambda}(v)$ coming from $v\in X^{*}$ define the mixed density
$x=\rho-\varepsilon J^{-1}(v).$ (57)
This definition is unique only if a Hohenberg–Kohn-type result is available so that the corresponding $v\in-\partial F^{\lambda}(\rho)$ is unique, but even without uniqueness we can make this assignment and will simply allow multiple such combinations. Choosing $\varepsilon>0$ small means that in order to get the generalized density, we only mix in a small quantity of the corresponding potential $v$. We note that the resulting $x$ does not need to be positive, normalizable, or have a finite kinetic energy, only the $v$-representable density $\rho$ has all those properties. With $v\in X^{*}$ as an external potential with assigned density $\varepsilon J^{-1}(v)$ that gets added to the internal electronic density of a quantum system, one could call $x$ a ‘bidensity’. It has been given different names before, such as ‘depot’ (for density–potential), ‘pseudodensity’ [1], or ‘quasidensity’ [54]. We will not commit to a singular nomenclature here, and instead refer to it as a general element $x\in X$ or a ‘mixed density’ when its decomposition like in Eq. (57) matters.
One physically motivated possibility is to choose the spaces such that $J$ corresponds to solving the Poisson equation of electrostatics and $\varepsilon J^{-1}(v)=-\varepsilon\Delta v$. This choice was previously suggested in the context of density–potential inversion [4]; it will be covered in more detail in Section VI.1 and has already found multiple use in ameliorating the Kohn–Sham method [55; 56; 57], albeit not with the connection to the topology of density and potential space.
Since $\rho\in\partial E^{\lambda}(v)=\partial(F^{\lambda})^{*}(-v)$ from Eq. (32) (not forgetting a minus from the chain rule) we note that remarkably, Eq. (57) reproduces Eq. (18) from the properties of the MY regularization. We just have to use $E_{\varepsilon}^{\lambda}(v)=-(F_{\varepsilon}^{\lambda})^{*}(-v)$ from Eq. (50) and the homogeneity of the duality map to get
$x\in\partial E_{\varepsilon}^{\lambda}(v)=\partial E^{\lambda}(v)-\varepsilon J^{-1}(v),$ (58)
which is just the superdifferential of Eq. (53). Since $\partial E_{\varepsilon}^{\lambda}$ has a well-defined inverse, $\nabla F_{\varepsilon}^{\lambda}$, it follows that for any $x\in X$ we can find a $v\in X^{*}$ such that the representation in Eq. (57) holds. In other words, every $x\in X$ can be represented as a mixed density. The regularizing effect of the mixing for $F^{\lambda}(\rho)$ is now easily explained: If $v\neq v^{\prime}$ lead to (almost) the same ground-state densities $\rho\approx\rho^{\prime}$, then this will not be true any more for $x=\rho-\varepsilon J^{-1}(v)$ and $x^{\prime}=\rho^{\prime}-\varepsilon J^{-1}(v^{\prime})$ due to the additional external term.
Also the proximal map receives a simple interpretation. If we set $v=-\nabla F_{\varepsilon}^{\lambda}(x)$ in Eq. (57) and solve for $\rho$, it follows from property v in Section II.3 that
$\rho=x-\varepsilon J^{-1}(\nabla F_{\varepsilon}^{\lambda}(x))=\mathop{\mathrm{prox}}_{\varepsilon F^{\lambda}}(x).$ (59)
Hence the internal density is just the proximal point of the mixed density. The consequence for DFT is that, to be able to work with the MY-regularized functionals, we also have to switch from internal densities $\rho\in D$ to general, mixed densities $x\in X$. The usual internal densities can then always be recovered as the proximal point or by performing the limit $\varepsilon\to 0$, where the limit procedure is well-defined through the properties of the MY regularization.
IV Density–potential inversion
The limit procedure in property vii of the MY regularization in Section II.3 that says $\lim_{\varepsilon\to 0}\nabla F_{\varepsilon}^{\lambda}(x)\in\partial F^{\lambda}(x)$ will be the basis for a method to calculate the corresponding external potential from a given ground-state density. The method itself is already known since a long time as the “ZMP method” [2], but our derivation as well as the accompanying interpretation is only feasible by resorting to MY regularization. We also add a more speculative subsection that makes an interesting and previously unexplored connection to the Dyson equation. In the section we will require Assumption 1 but sometimes, when it is mentioned, demand higher regularity of the involved spaces for additional results.
As a motivation, we start with the variational problem for the universal density functional from Eq. (37),
$F^{\lambda}(\rho)=\sup_{v\in X^{*}}\{E^{\lambda}(v)-\langle v,\rho\rangle\},$ (60)
whose stationarity condition $\partial E^{\lambda}(v)+\rho\ni 0$ gives $v\in-\partial F^{\lambda}(\rho)$ as an optimizer if $\rho$ is assumed $v$-representable. This approach to density–potential inversion was first applied to atoms by Colonna and Savin [58] in 1999 and by Teale, Coriani and Helgaker to molecules [59; 60; 61] in 2009. It sometimes called “Lieb optimization”, “Lieb maximization” [60; 61; 62] or the “dual problem” [63] in the literature. Convergence of this method can be facilitated by making the functional in the optimization strictly concave by subtracting $\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$, which corresponds to MY regularization,
$F_{\varepsilon}^{\lambda}(\rho)$ $=\sup_{v\in X^{*}}\{E_{\varepsilon}^{\lambda}(v)-\langle v,\rho\rangle\}$ (61)
$=\sup_{v\in X^{*}}\left\{E^{\lambda}(v)-\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}-\langle v,\rho\rangle\right\}.$
The stationarity condition for an optimizer $v_{\varepsilon}$ is then
$0\in\partial E^{\lambda}(v_{\varepsilon})-\varepsilon J^{-1}(v_{\varepsilon})-\rho.$ (62)
We use Eq. (58) to translate this to $\rho\in\partial E_{\varepsilon}^{\lambda}(v_{\varepsilon})$ which is equivalent to $v_{\varepsilon}=-\nabla F_{\varepsilon}^{\lambda}(\rho)$ by using the analogous relation to (39) for the regularized functionals. By property v from Section II.3, we can connect $v_{\varepsilon}$ to the proximal map,
$v_{\varepsilon}=\varepsilon^{-1}J(\mathop{\mathrm{prox}}_{\varepsilon F^{\lambda}}(\rho)-\rho).$ (63)
This equation then serves as the basis for MY density–potential inversion as an iterative procedure involving the proximal point of $\rho$, to find $v_{\varepsilon}$, which in the limit $\varepsilon\to 0$ gives the representing potential of $\rho$. The method will be more properly derived from MY regularization in the following section.
IV.1 Proximal-point iteration and relation to ZMP
The inverse problem of finding the corresponding potential $v$ for a given ground-state density $\rho$ is central to DFT and to achieve this, we ‘only’ have to evaluate the subdifferential of $F^{\lambda}(\rho)$ (with a minus sign, see Eq. (38)),
$v\in-\partial F^{\lambda}(\rho).$ (64)
As mentioned before, it is not guaranteed that this subdifferential is non-empty even for $\rho\in D\subsetneq X$ and, in non-standard settings for DFT (like DFT on finite lattices [32]), it may even be multi-valued. For the regularized functional, no such problems occur, and we can evaluate
$v_{\varepsilon}=-\nabla F_{\varepsilon}^{\lambda}(\rho),$ (65)
for any $\rho\in D$ and even for $\rho\in X\setminus D$. Of course, this does in general not yield the same potential as Eq. (64), but it does if $\rho=\mathop{\mathrm{prox}}_{\varepsilon F^{\lambda}}(\rho)$ by property v in Section II.3 or as $\varepsilon\to 0$ by property vii. We therefore know how the functional derivative in Eq. (65) relates back to the non-regularized form, but we may wonder how it can be computed.
Noting relation v for the proximal map and introducing $p=\mathop{\mathrm{prox}}_{\varepsilon F^{\lambda}}(\rho)$, we obtain
$\nabla F_{\varepsilon}^{\lambda}(\rho)=\varepsilon^{-1}J(\rho-p)\in\partial F(p).$ (66)
which by Eq. (38) gives
$p\in\partial E^{\lambda}(\varepsilon^{-1}J(p-\rho)).$ (67)
Now we have the unknown $p\in D$ on both sides, so this form lends itself to an iteration scheme. Beginning with a guess $p_{1}$, we iterate
$p_{i+1}\in\partial E^{\lambda}(\varepsilon^{-1}J(p_{i}-\rho)).$ (68)
Due to the possibility of degeneracy in the ground state, the next step $p_{i+1}$ is possible non-unique, which creates a possible issue that has not been explored up to date. If converged, this process yields the proximal point $p=\mathop{\mathrm{prox}}_{\varepsilon F^{\lambda}}(\rho)$ and we retrieve Eq. (63),
$v_{\varepsilon}=-\nabla F_{\varepsilon}^{\lambda}(\rho)=\varepsilon^{-1}J(p-\rho),$ (69)
where the resulting potential $v_{\varepsilon}$ depends on the choice of the smoothing parameter $\varepsilon$. Repeating this iterative scheme for a sequence of decreasing $\varepsilon>0$, we can extrapolate to the limit $\varepsilon=0$ and obtain $v_{\varepsilon}\rightharpoonup v$. Under the minimal Assumption 1, convergence $v_{\varepsilon}\rightharpoonup v$ is weak; under the stronger Assumption 2, convergence is strong. By virtue of vii, the resulting potential $v$ is the element of $\partial F^{\lambda}(\rho)$ with the smallest norm.
The benefit of this method is that we do not need to evaluate the (usually unknown) $F^{\lambda}(\rho)$, relying instead on $\partial E^{\lambda}(v)$ by solving the Schrödinger equation. This density-potential inversion procedure using MY regularization was previously detailed in Penz, Csirik, and Laestadius [4].
IV.2 Moreau–Yosida in inverse Kohn–Sham
While the procedure from the preceding section remains a hard task for interacting particles, density–potential inversion is often considered for the non-interacting KS system ($\lambda=0$) and then typically specified as “inverse Kohn–Sham” DFT. Starting with a given ground-state density $\rho$ for some external potential $v$ in an interacting system that is also non-interacting $v$-representable, the Kohn–Sham potential must satisfy
$v_{\mathrm{s}}\in-\partial F^{0}(\rho).$ (70)
We can approach this potential according to Eq. (69) with
$v_{\mathrm{s}}=\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}J\left(\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho)-\rho\right),$ (71)
where the proximal map can be evaluated by using the iteration in Eq. (68). To obtain $\rho\in\partial E^{0}(v)$, we solve the non-interacting Schrödinger equation, introducing orbitals.
A more intuitive viewpoint for this reasoning is to compare the corresponding minimization of the MY regularization and the very KS idea. Since $\rho$ is by assumption a minimizer of $F^{0}(\cdot)+\langle v_{\mathrm{s}},\cdot\rangle$, and $\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho)$ is the minimizer of $F^{0}(\cdot)+\tfrac{1}{\varepsilon}\|\cdot-\rho\|^{2}$, we must have simultaneously
$0\in\partial F^{0}(\rho)\,\,$ $+\,\,$ $v_{\mathrm{s}},$ (72)
$0\in\partial F_{\varepsilon}^{0}(\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho))\,\,$ $+\,\,$ $\frac{1}{\varepsilon}J(\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho)-\rho).$ (73)
Motivated by the fact that $\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho)\to\rho$ as $\varepsilon\to 0$ one can then under technical assumptions prove that indeed $v_{\mathrm{s}}$ corresponds to the limit of $\tfrac{1}{\varepsilon}J(\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho)-\rho)$ as the equations above suggest.
Now, exactly this method was proposed and implemented as a density–potential inversion scheme by Zhao, Morrison, and Parr [2] (and was later termed “ZMP method”), where the choice of spaces corresponds to homogeneous Sobolev spaces for potentials and their duals for densities (also see Section VI.1 and an upcoming publication for periodic domains like in Section VII). But in the original publication the method was developed ad hoc and only proven to work empirically. Only quite recently, it was shown [4] that the proximal-point iteration with $\varepsilon\to 0$ relates closely to the ZMP method for the usual Sobolev spaces and that it also reproduces a method by van Leeuwen and Baerends [64] if employed on a Hilbert space with $J=J^{-1}=\mathop{\mathrm{id}}$. It is further equivalent to yet another density–potential inversion method by Kumar and Harbola [65], where different choices for their penalty function $S[\rho]$ would correspond to different topologies for $X$. The method has been analyzed numerically [19; 66] (see also Section VII here) and adds itself to a long list of KS inversion techniques [67]. More than that, it shows that MY regularization offers a mathematical framework that allows for an elegant unification of numerous density–potential inversion methods.
IV.3 Resolvent form and Dyson-like equation
Instead of establishing a practical scheme for density–potential inversion, this section will be devoted to an interesting reformulation of the basic relation $v=-\nabla F_{\varepsilon}^{\lambda}(x)$ that becomes possible for a regularized functional. Starting from Eq. (18) for the universal density functional $F^{\lambda}(x)$ and using Eq. (14) we get
$(\nabla F_{\varepsilon}^{\lambda})^{-1}(v)=\partial(F^{\lambda})^{*}(v)+\varepsilon J^{-1}(v).$ (74)
Even though $\nabla F_{\varepsilon}^{\lambda}(x)$ is single-valued, its inverse is in general a set-valued map, so the equation above is actually between sets. From Eq. (32) we get (not forgetting the minus from the chain rule) $\partial(F^{\lambda})^{*}(-v)=\partial E^{\lambda}(v)$ and we apply this to the function above after switching $v\mapsto-v$, where we also use that $J^{-1}$ is homogeneous. Then for $v=-\nabla F_{\varepsilon}^{\lambda}(x)$ (or, more directly, from Eq. (58)) we have
$(\nabla F_{\varepsilon}^{\lambda})^{-1}(-v)=\partial E^{\lambda}(v)-\varepsilon J^{-1}(v)=x.$ (75)
Finally, we invert the whole expression, arriving at a resolvent-type formula that can be applied to all elements from $X$, the density space, to yield the corresponding potential,
$\nabla F^{\lambda}_{\varepsilon}=(\partial E^{\lambda}-\varepsilon J^{-1})^{-1}.$ (76)
Note that in the Hilbert-space setting we already saw a resolvent form for the proximal map in property x in Section II.3. By transforming Eq. (75) into an iterative procedure to solve for $v$, we arrive at
$\nabla F_{\varepsilon}^{\lambda}=-\varepsilon^{-1}J+\varepsilon^{-1}J\circ\partial E^{\lambda}\circ(-\nabla F_{\varepsilon}).$ (77)
Here we used $\circ$’ to denote the concatenation of operations and, unsurprisingly, this result is equivalent to the proximal-point iteration in Eq. (68) if we combine it with Eq. (69). But writing it in the form above exposes a striking parallel to the Dyson equation [68], where $\nabla F_{\varepsilon}^{\lambda}$ takes the role of the interacting Green’s function (dressed propagator), $\varepsilon^{-1}J$ that of the free Green’s function (bare propagator), and $\partial E^{\lambda}$ (with an additional minus to get the sign right) would be the self-energy. Applying $\varepsilon J^{-1}$ from the left and $(-\nabla F_{\varepsilon}^{\lambda})^{-1}=\partial E_{\varepsilon}^{\lambda}$ from the right yields
$\partial E_{\varepsilon}^{\lambda}=\partial E^{\lambda}-\varepsilon J^{-1}$ (78)
which is just Eq. (58) again and shows another path to arrive at Eq. (77). In our setting of MY applied to DFT, the mapping $\varepsilon^{-1}J$, that gets defined by the topology of the space $X$ itself, gives the ‘bare’ connection between densities and potential. Compared to this, $\nabla F_{\varepsilon}^{\lambda}$ is a ‘dressed’ way how to connect the quantities, where the dressing occurs through the presence of particles the potential can interact with. Lastly, $\partial E^{\lambda}$ derives from the functional $E^{\lambda}(v)$, the energy of the matter system. This last element introduces a possible ambiguity into the formulation that needs to be considered if degeneracy occurs and $\partial E^{\lambda}(v)$ is a multi-valued set. Nevertheless, this reformulation allows for an interesting alternative viewpoint that links it to Green’s function methods and diagrammatic expansions.
V Regularized Kohn–Sham theory
We already noted in Section II.5 that the requirement of having all involved densities simultaneously interacting and non-interacting $v$-representable puts a severe limitation on the possibility of a rigorous formulation of the Kohn–Sham method. This limitation can also influence numerical applications of the method, since it means that basis-set limits are ill-defined. The MY regularized functional $F^{\lambda}_{\varepsilon}(x)$, with the interaction strength now again noted as $\lambda$, offers a way how to avoid representability issues since the functional is differentiable and every $x\in X$ becomes representable through $\nabla F^{\lambda}_{\varepsilon}(x)$. The corresponding regularized Kohn–Sham method was first formulated in Kvaal et al. [1] for a Hilbert-space setting and later extended to Banach spaces [14]. In this section, we not only give its basic formulation, but also show how usual exchange–correlation can be translated to the new method. Finally, we repeat the main steps of a convergence proof for the method with an additional damping step. As usual, we require Assumption 1 but have to move to the Hilbert-space setting (Assumption 3) and even finite-dimensional spaces for the convergence proof. Since in the KS method we ought to stay connected to the Schrödinger equation and related KS equation, also Assumption 4 ($F^{\lambda}=F_{\mathrm{DM}}^{\lambda}$) is necessary.
V.1 Iteration scheme
We will build directly on the iteration scheme introduced in Section II.5 but, with the help of MY regularization, can completely avoid the critical issues within the usual formulation of the KS method. As a reminder, we were able to introduce a well-defined exchange–correlation potential $v_{\mathrm{xc}}(\rho)$ only if $\rho$ is simultaneously interacting and non-interacting $v$-representable, i.e., at every density $\rho$ that occurs within the KS iteration both subdifferentials $\partial F^{1}(\rho)$ and $\partial F^{0}(\rho)$ must be non-empty. From what we learned about MY regularization, this can be directly guaranteed if instead we switch to the regularized functionals $F^{1}_{\varepsilon}(x)$ and $F^{0}_{\varepsilon}(x)$. However, we then need to make sense of the resulting regularized KS iteration procedure. So let us define in exact analogy to Eq. (40) for any $x\in X$ with $v=-\nabla F^{1}_{\varepsilon}(x)$ and $v_{\mathrm{s}}=-\nabla F^{0}_{\varepsilon}(x)$,
$v_{\mathrm{xc},\varepsilon}(x)$ $=v_{\mathrm{s}}-v-v_{\mathrm{H}}(x)$ (79)
$=\nabla F^{1}_{\varepsilon}(x)-\nabla F^{0}_{\varepsilon}(x)-v_{\mathrm{H}}(x)$
$=\nabla(\underbrace{F^{1}_{\varepsilon}(x)-F^{0}_{\varepsilon}(x)-E_{\rm H}(x)}_{\textstyle E_{\mathrm{xc},\varepsilon}(x)}).$
We now write $x$ instead of $\rho$ to highlight that this is an arbitrary element (see Section III.3) in $X$ that does not need to have the properties of an actual density. We can now define the same, self-consistent iteration procedure as in Eqs. (44) and (45), where we can use Eq. (58) (which is equivalent to Eq. (18)) for the superdifferential of the $E^{0}_{\varepsilon}$ functional. Let the external potential that acts on the fully interacting system be $v_{\mathrm{o}}+v$ (as before, the offset $v_{\mathrm{o}}$ will be included into the functionals), then iterate starting from an initial guess,
$v_{i+1}=v+v_{\mathrm{H}}(x_{i})+v_{\mathrm{xc},\varepsilon}(x_{i}),$ (80)
$x_{i+1}\in\partial E^{0}_{\varepsilon}(v_{i+1})=\partial E^{0}(v_{i+1})-\varepsilon J^{-1}(v_{i+1}).$ (81)
In solving for the next $x_{i+1}$ in Eq. (81), we rely on exactly the same non-interacting Schrödinger equation that allows an orbital construction as in the usual KS procedure, we just have to shift the resulting density by $\varepsilon J^{-1}(v_{i+1})$ as the density–potential mixing dictates. If the procedure converges (see Section V.3 for a convergence proof that applies if $X$ is finite-dimensional and an additional damping step is added) we get
$x_{i}\to x,\qquad v_{i}\to v_{\mathrm{s}},$ (82)
and the resulting $x$ can then be mapped to the ground-state density with the proximal map according to Eq. (59). But now an important difference to the usual KS procedure appears that was already apparent in the formulation of the regularized KS iteration: The method aims at a shared $x\in X$ for the full and auxiliary system, and not at the internal ground-state density. This means, when $v_{\mathrm{o}}+v_{\mathrm{s}}$ is applied to the auxiliary system we get the same $x=\rho_{\mathrm{s}}-\varepsilon J^{-1}(v_{\mathrm{s}})$ as when $v_{\mathrm{o}}+v$ is applied to the full system. Therefore, in general $\rho\neq\rho_{\mathrm{s}}$ and instead we have the ground-state densities
$\rho=x+\varepsilon J^{-1}(v)=\mathop{\mathrm{prox}}_{\varepsilon F^{1}}(x),$ (83)
$\rho_{\mathrm{s}}=x+\varepsilon J^{-1}(v_{\mathrm{s}})=\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(x).$ (84)
That it must hold $\rho,\rho_{\mathrm{s}}\in D$ serves as a consistency check for the method or can be used as a posteriori corrections. To get an expression for the ground-state energy, we can rely on the analogous expression to Eq. (48),
$E^{1}_{\varepsilon}(v)=E^{0}_{\varepsilon}(v)+E_{\mathrm{xc},\varepsilon}(x)-E_{\rm H}(x)-\langle v_{\mathrm{xc},\varepsilon}(x),x\rangle,$ (85)
that we just need to shift as in Eq. (17),
$E^{1}(v)=E^{1}_{\varepsilon}(v)+\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}.$ (86)
Note, however, that while we recover the usual ground-state energy $E^{1}(v)$ this way, it does not mean that the regularized Kohn–Sham method will yield the same orbitals as the usual method. This is clear since the orbitals in regularized Kohn–Sham will produce a density $\rho_{\mathrm{s}}$ that is different to the interacting density $\rho$, and instead the generalized densities $x=\rho-\varepsilon J^{-1}(v)=\rho_{\mathrm{s}}-\varepsilon J^{-1}(v_{\mathrm{s}})$ agree.
V.2 Transforming exchange–correlation potentials
While this reformulation of the KS iteration scheme with the help of MY regularization is conceptually already very appealing, the question remains how it relates back to the unregularized setting. More concretely, we want to know how $v_{\mathrm{xc},\varepsilon}(x)$ relates to $v_{\mathrm{xc}}(\rho)$, provided no $v$-representability issues arise at $\rho$. For this we take Eqs. (83) and (84) and combine them into
$v_{\mathrm{xc},\varepsilon}(x)=\varepsilon^{-1}J(\rho_{\mathrm{s}}-x)-\varepsilon^{-1}J(\rho-x)-v_{\mathrm{H}}(\rho).$ (87)
Note that if $X$ is a Hilbert space then $J$ is a linear mapping and the expression above can be simplified. Within the regularized KS iteration we already have access to $\rho_{\mathrm{s}}$ in every step as an element of $\partial E^{0}(v)$, so we just need the corresponding $\rho\in\partial E^{1}(v)$. Since we of course want to avoid solving the full problem, the $v_{\mathrm{xc}}(\rho)$ from standard KS-DFT (or any approximation to it) enters the picture, since it allows getting the same density from the auxiliary system by adding just the usual exchange–correlation potential, $\rho\in\partial E^{0}(v+v_{\mathrm{H}}(\rho)+v_{\mathrm{xc}}(\rho))$. Together with $v=\varepsilon^{-1}J(\rho-x)$, we can derive a self-consistent iteration that works similar to KS itself,
$\begin{array}[]{cccccccc}\rho&\in\partial E^{0}(\varepsilon^{-1}J(&\rho&-x)+v_{\mathrm{H}}(&\rho&)+v_{\mathrm{xc}}(&\rho&))\\ \updownarrow&&\updownarrow&&\updownarrow&&\updownarrow&\\ \rho_{i+1}&\in\partial E^{0}(\varepsilon^{-1}J(&\rho_{i}&-x)+v_{\mathrm{H}}(&\rho_{i}&)+v_{\mathrm{xc}}(&\rho_{i}&)).\end{array}$ (88)
If this iteration is integrated into the regularized KS scheme from Section V.1 together with determining $v_{\mathrm{xc},\varepsilon}(x)$ from Eq. (87), we can recycle any given exchange–correlation approximation for the regularized setting. It must be added that this transformation of exchange–correlation potentials for the regularized KS method was not implemented to date.
V.3 Convergence proof
Now that a rigorous formulation of the KS method has been achieved, the question of convergence of the method to the desired ground-state density, when using the exact exchange–correlation functional, remains to be answered. First results on convergence of the usual KS method (Section II.5) required a $v$-representability assumption and only showed convergence in energy [69; 70]. With the help of MY regularization the problem of $v$-representability could be avoided and by the additional inclusion of a damping step a full convergence proof was achieved at least for a setting where $X$ is a (finite-dimensional) Euclidean space [71; 72; 20]. This setting means that Assumption 3 holds, and we slightly generalize the result to any finite-dimensional Hilbert space $X$ in the following presentation.
Implementations of KS regularly come with additional steps like damping that facilitate convergence. Such a damping step must also be included here, which means that in Eq. (81) not the full step from $x_{i}$ to $x_{i+1}$, but one with a reduced step length is applied. This step length will be such that one reaches exactly the apex of the regularization parabola that is tangentially aligned to $F^{1}_{\varepsilon}+v$ at $x_{i}$, as illustrated in Figure 2. Yet, it must be stressed that such a choice of step-length is not practical, since it requires knowledge about the correct functional. Its sole purpose here is to be able to establish a mathematical convergence proof. Different schemes with an a priori or adaptive choice of step length can be constructed as well [50; 14; 73]. We also note on a strong parallel to the proximal-point algorithm [74], where the minimum of $f(x)$ is found by iterating $x_{i+1}=\mathop{\mathrm{prox}}_{\varepsilon f}(x_{i})$. By the definition of the proximal map, Eq. (16), this corresponds to jumping to the apex of the full regularization function $\varepsilon\phi(x-x_{i})$ attached to $x_{i}$ in every step. The important difference to the regularized Kohn–Sham method is that the method fixes the step direction in the previous step (by solving the non-interacting system with the exchange–correlation potential) and then jumps to the apex of the regularization parabola restricted to this direction.
[…]
Figure 2: Illustration of one iteration step in the regularized KS scheme.
The regularized KS scheme with additional damping step is then given by the iteration
$v_{i+1}=v+v_{\mathrm{H}}(x_{i})+v_{\mathrm{xc},\varepsilon}(x_{i}),$ (89)
$x_{i+1}^{\prime}\in\partial E^{0}_{\varepsilon}(v_{i+1})=\partial E^{0}(v_{i+1})-\varepsilon J^{-1}(v_{i+1})$ (90)
$\quad\text{and set}\quad y_{i}=(x_{i+1}^{\prime}-x_{i})/\|x_{i+1}^{\prime}-x_{i}\|_{X},$
$x_{i+1}=x_{i}+\tau_{i}y_{i},$ (91)
$\quad\text{where}\quad\tau_{i}=-\varepsilon\langle\nabla F^{1}_{\varepsilon}(x_{i})+v,y_{i}\rangle.$
In the following, we provide most steps and explanations from the original proof [71] that was subsequently revised due to an error in the first version [72]. For additional details we refer to the original publications.
From the first step and using the definition of $v_{\mathrm{xc},\varepsilon}$ in Eq. (79) we have that
$\nabla F^{1}_{\varepsilon}(x_{i})+v=v_{i+1}+\nabla F^{0}_{\varepsilon}(x_{i}).$ (92)
The iteration is converged if either side of this equation is zero, which by the inverse subdifferential relation, Eq. (14), means that $x_{i}\in\partial E^{1}_{\varepsilon}(v)$ and also $x_{i}\in\partial E^{0}_{\varepsilon}(v_{i+1})$ (the interacting and the non-interacting system share the same $x$ as in Eqs. (83) and (84)). In order to track progression of the method, we define
$e_{i}=F_{\varepsilon}^{1}(x_{i})+\langle v,x_{i}\rangle.$ (93)
In case of correct convergence $x_{i}\to x\in\partial E^{1}_{\varepsilon}(v)$ we then have $\lim e_{i}=E^{1}_{\varepsilon}(v)$, which equals the ground-state energy of the interacting system minus $\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$. We take the expression of Eq. (93) that is differentiable and convex in $x_{i}$ and rewrite its directional derivative in the chosen direction $x_{i+1}^{\prime}-x_{i}$ with the help of Eq. (92),
$\langle\nabla F^{1}_{\varepsilon}(x_{i})+v,x_{i+1}^{\prime}-x_{i}\rangle=\langle v_{i+1}+\nabla F^{0}_{\varepsilon}(x_{i}),x_{i+1}^{\prime}-x_{i}\rangle.$ (94)
Notice that $x_{i+1}^{\prime}\in\partial E^{0}_{\varepsilon}(v_{i+1})$ and $x_{i}\in\partial E^{0}_{\varepsilon}(-\nabla F^{0}_{\varepsilon}(x_{i}))$ by definition, so from Section II.3 we can use the strong monotonicity property xii of $\partial E^{0}_{\varepsilon}(v)=\partial(F^{0}_{\varepsilon})^{*}(-v)$ (note the change of sign) to get
$\langle\nabla F^{1}_{\varepsilon}(x_{i})+v,x_{i+1}^{\prime}-x_{i}\rangle$ $\leq-\varepsilon\|v_{i+1}+\nabla F^{0}_{\varepsilon}(x_{i})\|^{2}_{X^{*}}$ (95)
$=-\varepsilon\|\nabla F^{1}_{\varepsilon}(x_{i})+v\|^{2}_{X^{*}}.$
This means that by a step in the direction $x_{i+1}^{\prime}-x_{i}$, which is parallel to $y_{i}$, we can always decrease the energy, if the step length is not taken too large. For this reason we limit the step length to $\tau_{i}$ such that we just reach the apex of the regularization parabola (at value $m_{i}$). We get this value from equating the directional derivatives of $F^{1}_{\varepsilon}(x_{i})+v$ and the regularization parabola at $x_{i}$, which yields
$\langle\nabla F^{1}_{\varepsilon}(x_{i})+v,y_{i}\rangle=-\frac{1}{\varepsilon}\|x_{i+1}-x_{i}\|_{X}=-\frac{\tau_{i}}{\varepsilon}$ (96)
and guarantees $e_{i+1}\leq m_{i}<e_{i}$. Since the energy is bounded below and thus strictly decreasing, we already showed convergence in energy. By determining $e_{i}-m_{i}$ from the regularization parabola and using Eq. (96) we get
$\frac{\tau_{i}^{2}}{2\varepsilon}=\frac{1}{2\varepsilon}\|x_{i+1}-x_{i}\|_{X}^{2}=e_{i}-m_{i}\leq e_{i}-e_{i+1}\to 0$ (97)
that shows $\tau_{i}\to 0$. Taking Eqs. (95) and (96) together gives the estimate
$\frac{\tau_{i}}{\varepsilon}\|x_{i+1}^{\prime}-x_{i}\|_{X}\geq\varepsilon\|\nabla F^{1}_{\varepsilon}(x_{i})+v\|^{2}_{X^{*}}.$ (98)
If we can be sure that $\|x_{i+1}^{\prime}-x_{i}\|_{X}$ is bounded this means that $\nabla F^{1}_{\varepsilon}(x_{i})\to v$.
In order to show convergence in densities and potentials, a compactness argument must be used. The revised proof [72] shows boundedness of the sequence $(x_{i})_{i}$ simply from the normalization condition in a finite-dimensional Hilbert space $X$. This means $(x_{i})_{i}$ has (possibly many) convergent subsequences of which we exemplary take $x_{\alpha(i)}\to z$. Since we are in a Hilbert-space setting (Assumption 3), property ix holds and $\nabla F^{\lambda}_{\varepsilon}$ is Lipschitz continuous. Now the first step Eq. (89) asserts
$v_{\alpha(i)+1}$ $=v+v_{\mathrm{H}}(x_{\alpha(i)})+v_{\mathrm{xc},\varepsilon}(x_{\alpha(i)})$ (99)
$=v+\nabla F^{1}_{\varepsilon}(x_{\alpha(i)})-\nabla F^{0}_{\varepsilon}(x_{\alpha(i)}),$
so by continuity $v_{\alpha(i)+1}\to v+\nabla F^{1}_{\varepsilon}(z)-\nabla F^{0}_{\varepsilon}(z)$ and the subsequence is bounded. By the same argument as before also $x_{\alpha(i)+1}^{\prime}\in\partial E^{0}_{\varepsilon}(v_{\alpha(i)+1})$ is bounded and so we really have $\lim_{i\to\infty}(-\nabla F^{1}_{\varepsilon}(x_{\alpha(i)}))=-\nabla F^{1}_{\varepsilon}(z)=v$ from Eq. (98). This already implies $\lim_{i\to\infty}v_{\alpha(i)+1}=-\nabla F^{0}_{\varepsilon}(z)$, so we have found a potential for the non-interacting KS system reproducing $z\in\partial E^{1}_{\varepsilon}(v)$. Taking a different subsequence $x_{\beta(i)}$ can then yield another element from $\partial E^{1}_{\varepsilon}(v)$ as well as its corresponding KS potential.
VI Auto-regularization in mean-field theories
In certain, specialized settings, like the Hartree-approximation to DFT or DFT with induced magnetic fields (Maxwell–Schrödinger DFT), the special form of the additional energy terms leads to automatic MY regularization. Interestingly, in the first case this is not on the side of the universal density functional, as in Eq. (49), but affects the ground-state energy $E^{\lambda}(v)$. In the case of Maxwell–Schrödinger DFT we can apply Legendre–Fenchel transformations to two different arguments and thus have the possibility to define four different functionals from convex conjugation, in some analogy to the potentials of thermodynamics that are also linked by Legendre transformations (the one-dimensional version of a Legendre–Fenchel transform). There, it turns out, that two of those functionals can even be linked to a certain form of double MY regularization called Lasry–Lions regularization.
VI.1 Hartree auto-regularization
Through a special choice for the potential space previously mentioned in Sections III.2 and III.3, we can achieve automatic regularization of the energy functional in the Hartree approximation of DFT. This choice is the homogeneous Sobolev space [75; 76],
$X^{*}=\{u\in L^{2}(\mathbb{R}^{3})\mid\nabla u\in L^{2}(\mathbb{R}^{3}:\mathbb{R}^{3})\},$ (100)
that must be taken modulo constants in order to make $\|u\|_{X^{*}}=\sqrt{\varepsilon_{0}}\|\nabla u\|_{L^{2}}$ a valid norm. This includes the customary “up to an additive constant” already into the definition of the potential space. Note that we also built the vacuum permittivity $\varepsilon_{0}$ ($1/(4\pi)$ in atomic units) into the definition of the norm for reasons that will soon become evident. This space is actually a Hilbert space with inner product $\langle u,v\rangle_{X^{*}}=\varepsilon_{0}\langle\nabla u,\nabla v\rangle_{L^{2}}$, so we have Assumption 3 fulfilled. Applying partial integration in the $L^{2}$ inner product gives
$\langle u,v\rangle_{X^{*}}=-\varepsilon_{0}\int u(\bm{r})\Delta v(\bm{r})\mathrm{d}\bm{r}=\langle u,J^{-1}(v)\rangle,$ (101)
where the last term stems from the Riesz representation Eq. (7) applied to $X^{*}$ and means that $u\in X^{*}$ is applied to $J^{-1}(v)\in X$. But this allows us to write the inverse duality map as
$J^{-1}(v)=-\varepsilon_{0}\Delta v,$ (102)
so evaluating $v=J(x)$ is equivalent to solving the Poisson equation $\varepsilon_{0}\Delta v=x\in X$. Strictly speaking, the integral in Eq. (101) makes only sense if $\Delta v$ is still an $L^{2}$ function, but we can still take Eq. (102) in the sense of distributions. This implies that a general element $x\in X$ can be a distribution. Since $J$ is the canonical mapping between densities and potentials, and is itself fully defined by the topology (norm) of the space, we needed to include $\varepsilon_{0}$ in the norm’s definition already to get the units right (although in atomic units one would not notice). Now for regular distributions $x\in X$ (that can be represented by an $f\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{3})$ as $\langle u,x\rangle=\int u(\bm{r})f(\bm{r})\mathrm{d}\bm{r}$) we know the solution of $\varepsilon_{0}\Delta v=x$ to be exactly the Hartree potential (in this context also called the Riesz potential; note again that in atomic units $4\pi\varepsilon_{0}=1$),
$v(\bm{r})=\int\frac{f(\bm{r}^{\prime})}{|\bm{r}-\bm{r}^{\prime}|}\mathrm{d}\bm{r}^{\prime}=v_{\mathrm{H}}(f)(\bm{r}).$ (103)
Thus we have found an explicit expression for the duality map, $J(x)=v_{\mathrm{H}}(f)$, when $x\in X$ is a regular distribution represented by $f\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{3})$. Note that we usually use the symbols for distributional densities ‘$x$’ and their representing functions ‘$f$’ for elements of $X$ interchangeably, so we could have also written $J(x)=v_{\mathrm{H}}(x)$. Now take the norm in $X$ with one-half in front,
$\frac{1}{2}\|v\|_{X^{*}}^{2}=\frac{\varepsilon_{0}}{2}\|\nabla v\|_{L^{2}}^{2}=\frac{\varepsilon_{0}}{2}\int|\nabla v(\bm{r})|^{2}\mathrm{d}\bm{r},$ (104)
and we can recognize this as the electrostatic field energy. We can equivalently rewrite the norm in $X$ with the dual norm as in Eq. (5),
$\frac{1}{2}\|v\|_{X^{*}}^{2}=\frac{1}{2}\|x\|_{X}^{2}=\frac{1}{2}\langle J(x),x\rangle$ (105)
$=\frac{1}{2}\iint\frac{f(\bm{r})f(\bm{r}^{\prime})}{|\bm{r}-\bm{r}^{\prime}|}\mathrm{d}\bm{r}\mathrm{d}\bm{r}^{\prime}=E_{\rm H}(f),$
which retrieves the Hartree mean-field energy from Eq. (42) that is also the energy of the field induced by the charge distribution $f\in X$. Through these identifications, it becomes apparent that this choice of density and potential space is a very natural one for DFT and we believe that it can have a future importance for the formulation of DFT. Indeed, the use of these spaces in DFT was already suggested around 2012, based on an appearance in regularization of optimized-effective-potential methods [77], but never published. Still, several important mathematical questions are left to be studied in this setting, especially the self-adjointness and boundedness below of the Hamiltonian and lower semicontinuity of the ensemble constrained-search functional (Assumption 4). It is expected that the last point only holds on a bounded domain or through including a confining offset potential $v_{\mathrm{o}}$.
A very basic approximation for an interacting system is the Hartree functional
$F^{\rm Hartree}(\rho)=F^{0}(\rho)+E_{\rm H}(\rho)=F^{0}(\rho)+\frac{1}{2}\|\rho\|_{X}^{2},$ (106)
which becomes strictly convex through the addition of the Hartree mean-field energy. We are now prepared to show the auto-regularization of the corresponding energy functional from the Legendre–Fenchel transform (the attentive reader might already have noticed that a strictly convex functional $F^{\rm Hartree}(\rho)$ implies a MY-regularized convex conjugate functional). With $v=J(x)$ we get
$E^{\rm Hartree}(v)=-(F^{\rm Hartree})^{*}(-v)$ (107)
$=\inf_{\rho}\left\{F^{\rm Hartree}(\rho)+\langle v,\rho\rangle\right\}$
$=\inf_{\rho}\left\{F^{0}(\rho)+\frac{1}{2}\|\rho\|_{X}^{2}+\langle v,\rho\rangle\right\}$
$=\inf_{\rho}\left\{F^{0}(\rho)+\frac{1}{2}\|\rho\|_{X}^{2}+\langle v,\rho\rangle+\frac{1}{2}\|v\|_{X^{*}}^{2}\right\}-\frac{1}{2}\|v\|_{X^{*}}^{2}$
$=\inf_{\rho}\left\{F^{0}(\rho)+\frac{1}{2}\|\rho+x\|_{X}^{2}\right\}-\frac{1}{2}\|v\|_{X^{*}}^{2}$
$=F^{0}_{\varepsilon=1}(-x)-\frac{1}{2}\|v\|_{X^{*}}^{2}=F^{0}_{\varepsilon=1}(-J^{-1}(v))-\frac{1}{2}\|v\|_{X^{*}}^{2}.$
We can note several things here. In the third line we recognize that the considered energy content of the system is the (non-interacting) kinetic energy $F^{0}(\rho)$ of the system, the potential energy from coupling to the external $v$, but also the energy of the induced field $E_{\rm H}(\rho)$. This sums up to give the self-interaction error of the Hartree approximation. The regularization parameter can be read off directly from the fifth line of Eq. (107) as $\varepsilon=1$. Without including the vacuum permittivity into the norm we can have $\varepsilon=\varepsilon_{0}$, but then the duality map does not directly map densities to potentials any more. The new argument of the regularized functional is $x=-J^{-1}(v)=\varepsilon_{0}\Delta v$ and can of course be negative as a general element in $X$. Note that now the energy functional $E^{\rm Hartree}(v)$ is the result of MY regularization, not the universal density functional as before. Even though we subtract $\frac{1}{2}\|v\|_{X^{*}}^{2}$, which reminds of the way MY regularization was introduced in Section III.2, the energy functional does not necessarily become strictly concave (yet it is concave) since $F^{0}_{\varepsilon}(-x)$ is convex. Since $\rho=\nabla E^{\rm Hartree}(v)$ is now unique, there arises no ambiguity in the ground-state density.
VI.2 Maxwell–Schrödinger DFT
Maxwell–Schrödinger DFT [78] is an extension of standard DFT with electromagnetic fields and internal currents that also considers induced magnetic fields. In this respect it is exactly equivalent to the Hartree method discussed before in Section VI.1, adding a current-current integral instead of the density-density integral of Eq. (105). Since the currents couple to an external vector potential, we include those into the description, just like in current DFT [42]. The vector potentials live in a special homogeneous and divergence-free (Coulomb gauged) Sobolev space that is similar to the previously mentioned homogeneous Sobolev space,
$Y^{*}=\{\in L^{2}(\mathbb{R}^{3},\mathbb{R}^{3})\mid\nabla\times\in L^{2}(\mathbb{R}^{3},\mathbb{R}^{3}),\nabla\cdot\}.$ (108)
This space must further be taken modulo constant vector fields in order to make $\|\|_{Y^{*}}=\mu_{0}^{-1/2}\|\nabla\times\|_{L^{2}}$ a valid norm. Here, we built the vacuum permeability $\mu_{0}$ ($4\pi/c^{2}$ in atomic units) into the norm of the space for vector potentials and it is again a Hilbert space. Since $\nabla\times$ describes a magnetic field, $\frac{1}{2}\|\|^{2}_{Y^{*}}=\frac{1}{2\mu_{0}}\|\|^{2}_{L^{2}}$ exactly corresponds to the magnetostatic field energy. The corresponding space for current densities is then its dual space $Y$, where the duality map $J:Y\to Y^{*}$ then means solving the Maxwell equation
$\Delta\nabla\times\mu_{0}\bm{j}.$ (109)
It was possible to rewrite this as a set of component-wise Poisson equations since in the application of the vector Laplacian $\Delta\nabla(\nabla\cdot)-\nabla\times(\nabla\times)$ the divergence term vanishes for $\in Y^{*}$. The current space $Y$ includes distributional vector fields and the norm for regular distributions $\bm{j}\in Y$ features a current-current integral,
$\frac{1}{2}\|\|_{Y^{*}}^{2}=\frac{1}{2}\|\bm{j}\|_{Y}^{2}=\frac{\mu_{0}}{2}\iint\frac{\bm{j}(\bm{r})\cdot\bm{j}(\bm{r}^{\prime})}{4\pi|\bm{r}-\bm{r}^{\prime}|}\mathrm{d}\bm{r}\mathrm{d}\bm{r}^{\prime}.$ (110)
After setting the functional-analytic stage, let us introduce the coupled Maxwell–Schrödinger system. We redefine the kinetic-energy operator so that it includes vector potentials $\in Y^{*}$ through minimal coupling,
$\hat{T}=-\frac{1}{2}\sum_{j=1}^{N}(-\mathrm{i}\nabla_{j}+)^{2}.$ (111)
In contrast to the original publication [78], we do not include spin degrees-of-freedom, since this makes the notation more economic and does not change the main argument about auto-regularization. We now define a Hamiltonian $\hat{H}^{\lambda=1}_{v,}$ like in Eq. (23) just replacing $\hat{T}$ with $\hat{T}$. We further always choose the fully interacting case with $\lambda=1$ here and will thus suppress this parameter in the notation. The ground-state energy for a potential pair $(v,)\in X^{*}\times Y^{*}$ is then
$E(v,)=\inf_{\Psi}\langle\hat{\Psi},H_{v,}\Psi\rangle,$ (112)
which can equivalently be defined with variation over density matrices. This shows that this functional is again concave in $v$, but has no special convexity property in , a fact related to diamagnetism [78]. We define the corresponding constrained-search functional over ensembles [79; 80; 81],
[…]
that is now defined on $X\times Y$, the space of pairs of densities and paramagnetic currents. The notation $\hat{\Gamma}\mapsto(\rho,\bm{j}^{\mathrm{p}})$ means that in the variation all density matrices that map to a given one-particle density $\rho$ and paramagnetic current $\bm{j}^{\mathrm{p}}$ are considered. We then have
$E(v,)=\inf_{\begin{subarray}{c}\rho\in X\\ \bm{j}^{\mathrm{p}}\in Y\end{subarray}}\left\{F(\rho,\bm{j}^{\mathrm{p}})+\langle v+\tfrac{1}{2}^{2},\rho\rangle+\langle\bm{j}^{\mathrm{p}}\rangle\right\}.$ (114)
One immediately sees that the density now additionally couples to $\tfrac{1}{2}^{2}$ which means the resulting energy functional is also not concave in . This can easily be repaired by substituting $u=v+\tfrac{1}{2}^{2}$ which also makes the energy functional jointly concave in $u$ and . Yet note that this construction requires $^{2}\in X^{*}$, the space of potentials, which only holds when the spaces $X$ and $Y$ have a special compatibility property [15].
Without requiring joint concavity, the energy functional for Maxwell–Schrödinger DFT is defined as
$E_{\rm M}(v,\bm{A})=\inf_{\bm{\alpha}\in Y^{*}}\left\{E(v,\bm{\alpha}+\bm{A})+\frac{1}{2}\|\bm{\alpha}\|_{Y^{*}}^{2}\right\}.$ (115)
Here and in the following, we exactly follow the notation of Tellgren [78] in order to allow for direct comparison. The stationarity condition for this energy functional then gives the Schrödinger equation with Hamiltonian $\hat{H}_{v,\bm{\alpha}+\bm{A}}$ coupled to the Poisson equations for all components of the induced field $\bm{\alpha}\in Y^{*}$,
$\Delta\bm{\alpha}=\mu_{0}J^{-1}(\bm{\alpha})=\mu_{0}\bm{j}_{\Psi}.$ (116)
Here, $\bm{j}_{\Psi}$ is the physical current from the solution of the Schrödinger equation. Next, we define ${\bar{E}}_{\rm M}$ by subtracting the field energy $\frac{1}{2}\|\bm{A}\|^{2}_{Y^{*}}$. Also substituting $\bm{\alpha}+\bm{A}$, we see that we have arrived at a formula in close analogy to the Hartree auto-regularization in Eq. (107) (line 5), just for the vector potential,
${\bar{E}}_{\rm M}(v,\bm{A})=E_{\rm M}(v,\bm{A})-\frac{1}{2}\|\bm{A}\|^{2}_{Y^{*}}$ (117)
$=\inf_{\in Y^{*}}\left\{E(v,)+\frac{1}{2}\|\bm{A}\|_{Y^{*}}^{2}\right\}-\frac{1}{2}\|\bm{A}\|^{2}_{Y^{*}}.$
This functional is now jointly concave in $v$ and $\bm{A}$, yet we do not have the same effect of auto-regularization as for $E^{\rm Hartree}$ since $E(v,)$ is not convex (nor concave) in . For a regularized functional a further step is necessary.
By Legendre–Fenchel transform in just one or both arguments, one can construct three further functionals,
${\bar{F}}_{\rm M}(\rho,\bm{A})=\sup_{v\in X^{*}}\{{\bar{E}}_{\rm M}(v,\bm{A})-\langle v,\rho\rangle\},$ (118)
${\bar{e}}_{\rm M}(v,\bm{j}^{\mathrm{p}})=\sup_{\bm{A}\in Y^{*}}\{{\bar{E}}_{\rm M}(v,\bm{A})-\chi\langle\bm{A},\bm{j}^{\mathrm{p}}\rangle\},$ (119)
${\bar{f}}_{\rm M}(\rho,\bm{j}^{\mathrm{p}})=\sup_{\bm{A}\in Y^{*}}\{{\bar{F}}_{\rm M}(\rho,\bm{A})-\chi\langle\bm{A},\bm{j}^{\mathrm{p}}\rangle\}.$ (120)
In this step we introduced a new magnetic coupling constant $\chi$. Now it turns out that the last two functionals, ${\bar{e}}_{\rm M}(v,\bm{j}^{\mathrm{p}})$ and ${\bar{f}}_{\rm M}(\rho,\bm{j}^{\mathrm{p}})$, can be given in the form of Lasry–Lions regularizations [82], a form of double MY regularization, and that by a theorem due to Attouch and Aze [83] they both become Fréchet differentiable if $\chi>1$. This differentiability allows for a unique mapping from paramagnetic currents to vector potentials $\bm{A}$, either for given $v$ or $\rho$, and thus a form of a Hohenberg–Kohn theorem in the context of Maxwell–Schrödinger DFT. The details of this procedure can be found in the original publication by Tellgren [78].
If we remember that the Coulomb potential arises from a QED treatment of the longitudinal component of the electric field in Coulomb gauge [84], then the relation between the Hartree approximation (with only mean-field interaction) and standard DFT is the same as between Maxwell–Schrödinger DFT and a full (non-relativistic) QED treatment in terms of a functional theory. The development of such a theory, called QEDFT, is an ongoing program [85; 86; 87; 24; 88] with important implications for fields like polaritonic chemistry [89], and it might greatly benefit from MY regularization on its theoretical side.
VII Application of MY-based Kohn–Sham inversion to periodic settings
Crucial for the density–potential inversion described in Section IV is the computation of the proximal point $\mathop{\mathrm{prox}}_{\varepsilon F^{\lambda}}(x)$. For the full universal density functional, including interactions, this is a daunting task. Considering the inversion for the non-interacting functional $F^{0}(\rho)$ is considerably easier and the resulting method is called “inverse KS” since it relates to the KS system. For a periodic setting a KS inversion algorithm based on MY regularization has recently been implemented and analyzed leading to first error bounds [19].
A key advantage of the periodic setting for exploration is that natural choices for the density and potential spaces, $X$ and $X^{*}$, are the periodic Sobolev spaces $H_{\mathrm{per}}^{s}(\Omega)$ with $\Omega\subset\mathbb{R}^{3}$ denoting the periodically repeated unit cell [16, Sec. 2]. These spaces are Hilbert spaces (thus fulfill Assumption 3) and are easily characterized mathematically using Fourier modes $e_{\bm{G}}=e^{\mathrm{i}\bm{G}\cdot\bm{r}}/\sqrt{|\Omega|}$, namely by
$H_{\mathrm{per}}^{s}(\Omega,\mathbb{C}):=\left\{u=\sum_{\bm{G}}\hat{u}_{\bm{G}}e_{\bm{G}}\middle|\left\|u\right\|_{H_{\mathrm{per}}^{s}}<\infty\right\}$ (121)
with
$\left\|u\right\|_{H_{\mathrm{per}}^{s}}^{2}=\sum_{\bm{G}}(1+|\bm{G}|^{2})^{s}|\hat{u}_{\bm{G}}|^{2},$ (122)
and where the sum is over the usual reciprocal lattice vectors $\bm{G}$. This characterization is very much in line with the usual discretization schemes based on plane waves, where essentially one just selects a subspace of $H_{\mathrm{per}}^{s}(\Omega)$ by limiting the largest $|\bm{G}|$. Note that the space with $s=+1$ is different to the homogeneous Sobolev space of Eq. (100), since the $1+|\bm{G}|^{2}$ (instead of just $|\bm{G}|^{2}$) in the definition of the norm in Eq. (122) takes care of lower-order derivatives as well.
As a result of this definition, key quantities of the MY program, such as the duality map $J$, can be obtained in computable, closed-form expressions. For example, with the particular choice $X=H_{\mathrm{per}}^{-1}(\Omega)$ for densities and $X^{*}=H_{\mathrm{per}}^{1}(\Omega)$ for potentials one obtains the duality map in the Fourier representation as [19]
$J(\rho)=\sum_{\bm{G}}\frac{\hat{\rho}_{\bm{G}}e_{\bm{G}}}{1+\left|\bm{G}\right|^{2}}.$ (123)
Noticing that $1/(\mu+\left|\bm{G}\right|^{2})$ is the Fourier representation of the Yukawa potential $\Phi^{\mu}(\bm{r})=\exp{(-\mu\left|\bm{r}\right|)}/(4\pi\left|\bm{r}\right|)$ then admits the closed-form formulation as a convolution,
$J(\rho)(\bm{r})=(\Phi^{\mu=1}\ast\rho)(\bm{r})=\int_{\mathbb{R}^{3}}\frac{e^{-|\bm{r}-\bm{r}^{\prime}|}\rho(\bm{r}^{\prime})}{4\pi|\bm{r}-\bm{r}^{\prime}|}\mathrm{d}\bm{r}^{\prime},$ (124)
which is efficiently computable in a plane-wave discretization using Eq. (123) and fast-Fourier transforms. We remark that the validity of Assumption 4 must still explicitly be demonstrated with this choice of spaces, although the periodic setting naturally guarantees the existence of a ground state for every choice of potential $v\in H_{\mathrm{per}}^{1}(\Omega)$.
Targeting KS inversion in this setting we (as usual) assume that we are given a simultaneously interacting and non-interacting $v$-representable ground-state density $\rho_{\mathrm{gs}}$, meaning that there exists a KS potential $v_{\mathrm{s}}\in-\partial F^{0}(\rho_{\mathrm{gs}})$ and a potential $v\in-\partial F^{1}(\rho_{\mathrm{gs}})$. While $F^{0}(\rho)$ could be directly subjected to the MY program, results seem to generally improve if a maximal amount of known information about the problem is explicitly employed [2; 67]. We therefore choose a ‘guiding functional’
$\mathcal{F}(\rho)=F^{0}(\rho)+E_{\rm H}(\rho)+\langle v,\rho\rangle,$ (125)
which additionally includes the Hartree mean-field energy and the external potential. Noting that $E_{\rm H}(\rho)$ in the periodic setting is modified to include a compensating background, that is $E_{\rm H}(\rho)=\sum_{\bm{G}\neq 0}\left|\hat{\rho}_{\bm{G}}\right|^{2}/\left|\bm{G}\right|^{2}$, one shows $\mathcal{F}(\rho)$ to remain convex [19]. Lower semicontinuity of $F^{0}(\rho)$ and thus the entire functional $\mathcal{F}(\rho)$ in the $H_{\mathrm{per}}^{-1}$ topology remains to be shown. Similarly, we remark that the ideal choice of guiding functional is currently also an open research question.
Using property v of MY regularization in Section II.3 then yields
$\varepsilon^{-1}J(\rho_{\mathrm{gs,\varepsilon}}-\rho_{\mathrm{gs}})\in-\partial\mathcal{F}(\rho_{\mathrm{gs,\varepsilon}}),$ (126)
where we introduced the shorthand $\rho_{\mathrm{gs,\varepsilon}}=\mathop{\mathrm{prox}}_{\varepsilon\mathcal{F}}(\rho_{\mathrm{gs}})$. We remark that our assumptions on $\rho_{\mathrm{gs}}$ allow the definition of an exchange–correlation potential $v_{\mathrm{xc}}=v_{\rm s}-v-v_{\mathrm{H}}(\rho_{\mathrm{gs}})\in-\partial\mathcal{F}(\rho_{\mathrm{gs}})$ that we obtain in the limit $\varepsilon\to 0$ (just like in Section IV.1)
$v_{\mathrm{xc}}=\lim_{\varepsilon\to 0}v_{\mathrm{xc},\varepsilon}\quad\text{with}\quad v_{\mathrm{xc},\varepsilon}=\varepsilon^{-1}J(\rho_{\mathrm{gs,\varepsilon}}-\rho_{\mathrm{gs}}).$ (127)
To compute this numerically we require two ingredients:
  1. (i)
    an approach to compute the proximal points $\rho_{\mathrm{gs,\varepsilon}}$ and
  2. (ii)
    a suitable sequence of decreasing values for $\varepsilon$ enabling an extrapolation of the $\varepsilon\to 0$ limit.
One approach to achieve i is to follow the fixed-point iteration suggested in Eq. (68). However, in our previous work [19], we directly solve the minimization problem encoded in the proximal-point computation, i.e., we minimize
$\mathcal{F}(\rho)+\frac{1}{2\varepsilon}\|\rho-\rho_{\mathrm{gs}}\|_{H_{\mathrm{per}}^{-1}}^{2}$ (128)
over all $\rho\in D$, which automatically excludes non-physical densities (e.g., with infinite kinetic energies) from the minimization. Notably, for the numerical approximation of the supremum in Eq. (37) for $F^{0}(\rho)$ we employ the usual KS parametrization of the density in terms of orthonormal orbitals. This parametrization reads $\Phi=(\phi_{1},\ldots,\phi_{N})$ with $\langle\phi_{i},\phi_{j}\rangle=\delta_{ij}$ and occupations $\bm{f}=(f_{1},\ldots,f_{N})$ and gives
$\rho_{\Phi}(\bm{r})=\sum_{i}f_{i}|\phi_{i}(\bm{r})|^{2}.$ (129)
We can then write the proximal-point computation as the minimization of
$\mathcal{E}(\Phi,\bm{f})=$ $\,\sum_{i}f_{i}\|\nabla\phi_{i}\|^{2}_{L^{2}}+E_{\mathrm{H}}(\rho_{\Phi})$ (130)
$+\langle v,\rho_{\Phi}\rangle+\frac{1}{2\varepsilon}\|\rho_{\Phi}-\rho_{\mathrm{gs}}\|_{H_{\mathrm{per}}^{-1}}^{2}$
with respect to the orbitals $\phi_{i}$ and occupations $f_{i}$. In these expressions we do not make explicit any details about spin or Brillouin zone sampling, i.e., the choice of the $\bm{k}$-point grid to model periodic supercells [90, Ch. 2.8]. In that sense one could think of the index $i$ to run over all occupied orbitals of all spins and all reducible $\bm{k}$-points.
We remark that this energy expression Eq. (130) is closely related to the usual KS-DFT energy expression, where essentially the exchange–correlation energy term is just replaced by the penalty term of the regularization. Standard direct minimization techniques from KS-DFT are therefore directly applicable [91; 92; 93]. A first implementation of this approach in DFTK [19; 94] is capable to perform proximal-point computations for on non-trivial 3D solid-state systems, such as silicon, gallium arsenide or potassium chloride, providing thus a practical MY-based KS inversion procedure. Right now the treatment is still restricted to insulating solids, since the algorithm currently makes the assumption that all $f_{i}=1$ in Eq. (129), respectively only minimizes over the $N$ (spin) orbitals in Eq. (130). A generalization to $f_{i}\neq 1$, required for the treatment of metals, is possible taking inspiration from the KS-DFT literature [92; 95; 96; 97].
When minimizing the energy functional of usual KS problems, it is primarily the kinetic-energy term that deteriorates the conditioning due to its large-$\bm{G}$ divergence. Appropriate preconditioners are available to tame this behavior [98; 92]. While for large $\varepsilon$ values these are appropriate for the MY proximal point computation, for small $\varepsilon$ the penalty term $\frac{1}{2\varepsilon}\|\rho_{\Phi}-\rho_{\mathrm{gs}}\|_{H_{\mathrm{per}}^{-1}}^{2}$ can dominate. Refined preconditioning techniques for the MY setting that are $\varepsilon$-dependent and take both terms into account therefore need to be developed.
Figure 3: Convergence of $\rho$ and $v_{\mathrm{xc},\varepsilon}$ as a function of $\varepsilon$ for the bulk silicon test case and a basis truncation of $E_{\text{cut}}=30$, leading to an error $\Delta\rho=2.7\times 10^{-5}$. We refer the reader to the original reference [19] for details of the computational setup.
An additional complication in practical inverse KS calculations is that the input density $\rho_{\mathrm{gs}}$ itself is usually obtained from an inexact procedure, such as an experiment or a computation from a higher-level theory. As a result we do not actually have access to the exact $\rho_{\mathrm{gs}}$, but only a perturbed version $\tilde{\rho}_{\text{gs}}=\rho_{\mathrm{gs}}+\delta\rho$. Computing the proximal point for a given $\varepsilon$ with this perturbed ground state density thus only provides us with an estimate ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ instead of the correct $v_{\mathrm{xc},\varepsilon}$. Previous analysis [19] lets us deduce the bound
$\left\|v_{\mathrm{xc}}-{\tilde{v}}_{\mathrm{xc},\varepsilon}\right\|_{H_{\mathrm{per}}^{1}}$ $\leq\left\|v_{\mathrm{xc}}-v_{\mathrm{xc},\varepsilon}\right\|_{H_{\mathrm{per}}^{1}}+\left\|v_{\mathrm{xc},\varepsilon}-{\tilde{v}}_{\mathrm{xc},\varepsilon}\right\|_{H_{\mathrm{per}}^{1}}$ (131)
$\leq\left\|v_{\mathrm{xc}}-v_{\mathrm{xc},\varepsilon}\right\|_{H_{\mathrm{per}}^{1}}+\frac{C}{\varepsilon}\left\|\delta\rho\right\|_{H_{\mathrm{per}}^{-1}}$
for the error between the numerically accessible ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ and the desired true $v_{\mathrm{xc}}$. We note the two error contributions, where the first is controlled by taking $\varepsilon\to 0$. In contrast, the second term due to the inherent error $\delta\rho$ in the available ground-state density data will increase with decreasing $\varepsilon$. Therefore one usually obtains a convergence behavior similar to Figure 3 for the density and potential. Notably the potential $v_{\mathrm{xc},\varepsilon}$ initially converges towards the true $v_{\mathrm{xc}}$, but then diverges, suggesting that there is an optimal $\varepsilon$ value to employ to yield the most accurate estimate of the inverted potential.
The bound in Eq. (131) provides a qualitative explanation of the observed behavior as $\varepsilon\to 0$. However, it is neither sharp nor does it fully match the practical setting of periodic calculations, such that it is not yet sufficient to suggest an optimal $\varepsilon$ value a priori, let alone an entire $\varepsilon$ sequence for potential extrapolation. For example, for efficient plane-wave calculations employing a pseudopotential approximation is required. This implies the addition of a non-local (Kleinman–Bylander) term [99, Ch. 7] to the energy functional Eq. (130), which is outside of the scope of the current mathematical framework [19].
While the current analysis thus already provides first insights into the optimal choice of an $\varepsilon$ sequence (ingredient ii), it still requires considerable refinement until an a priori mathematically guided selection of such a sequence is feasible. An explicit numerical study of the $\varepsilon\to 0$ limit, e.g., based on a logarithmically spaced $\varepsilon$ grid, is thus generally required to empirically identify the optimal $\varepsilon$. However, the progress towards first analytical error bounds such as Eq. (131) as well as the rich literature on a priori and a posterior error analysis for the closely related problem of periodic KS-DFT [100; 101; 102; 103; 104; 105; 106; 107] make it hopeful that more refined and comprehensive analysis is feasible. A tantalizing outlook would be sharper error bounds, thus a better estimation of an optimal $\varepsilon$, or the combination of convergence properties of ${\tilde{v}}_{\mathrm{xc},\varepsilon}$ with extrapolation techniques to accelerate convergence.
Beyond refinements of the inversion algorithm itself, a number of possible choices in the mathematical setup have not yet been fully explored as well. This includes a different choice than $H_{\mathrm{per}}^{-1}(\Omega)$ for the density space $X$ and $H_{\mathrm{per}}^{1}(\Omega)$ for the potential space $X^{*}$. For example, the inverse choice ($H_{\mathrm{per}}^{-1}(\Omega)$ for the potential and $H_{\mathrm{per}}^{1}(\Omega)$ for the density) would in principle work as well and exactly fits to recent results about $v$-representability on periodic one-dimensional domains [17; 18]. Although it has to be noted that this will break the self-adjointness requirement from Assumption 4 in more than one dimension. We believe that systematically exploring the question of function spaces is a promising endeavor to precisely target the inversion procedure towards reproducing the physically relevant features of the potential: as visible from Eq. (130), the function space choice is directly reflected in the norms employed during the proximal point computation. From a physical point of view, the regions close to the nucleus are generally less relevant, such that an ideal function space $X$ should feature a norm that puts less weight on the core and larger weight on the valence regions.
In closing, we remark that the exceptionally close agreement between the plane-wave discretization basis and the Fourier characterization of the periodic Sobolev spaces make the periodic setting ideally suited for further joint numerical and analytical study of MY-regularized DFT. In combination with a flexible tool such as DFTK [94], mathematical ideas can be seamlessly implemented and tested both on toy problems as well as realistic materials systems. Considering specifically the similarity of MY-based KS inversion to other inversion approaches (such as the ZMP method), see Section IV, as well as the critical lack in mathematical analysis for practical inverse KS, this provides an accessible framework to close this important gap.
VIII Outlook
This work aimed at showing how, since its introduction into the field of DFT, the technique of MY regularization grew ever more useful and versatile. At this point, it is interesting to ask, which are the most promising routes to follow and what more can be expected. We here collect a number of ideas and possible research directions that we deem especially interesting and worthwhile to investigate. This will at the same time serve as a summary of the presented concepts relating to MY regularization applied to DFT.
  • In Section VI it was demonstrated how in certain mean-field approximations a choice of spaces for densities and potentials that is adapted such that $\|\cdot\|_{X^{*}}^{2}$ yields the field energy automatically leads to MY regularization. A very similar choice of spaces was made in the ZMP method for density–potential inversion (Section IV) and in the numerical discussion of the periodic setting (Section VII). In general, we believe that fitting the mathematical setting to physical principles, while taking all the functional-analytic considerations of Section II.4 into account, is extremely useful as a guiding principle. The original Lieb setting [3] with $X=L^{1}\cap L^{3}$ in some sense also had this in mind via the estimate of the $L^{3}$-norm of the density by the kinetic energy, but this is just comparative, while also an exact relation to the field energy is possible. Equation (102) demonstrated how this leads to an equivalence between the duality map and the electrostatic Poisson equation. It is thus possible to encode important physical relations fully into the geometry of the density and potential spaces. (Note that such an encoding of physical laws into geometry is also present quite generally in gauge theories through the introduction of fibre bundles.) This basic idea will be pursued further for a new formulation of QEDFT in a forthcoming publication.
  • The spaces of Section VI just mentioned also happen to be Hilbert spaces, thus fulfilling our Assumption 3. This means that additionally to the Banach-space structure several extremely beneficial properties hold true, like the linearity of the duality map and firm nonexpansiveness of the proximal map. Up to this point, this was not used to full capability. It is for example possible that the Hilbert-space setting can help to generalize the proof of convergence of the regularized KS scheme from Section V.3 to infinite dimensions.
  • A generalization in the opposite direction is also possible. This means considering Banach spaces and adapting the MY regularization itself to what is called the ‘modulus of convexity’ of the respective space. This is possible by replacing the $(\cdot)^{2}$ in Eq. (15) by $(\cdot)^{p}$, $p>1$, or right away by a Young function $\phi(\cdot)$. One then also needs to switch to a corresponding generalized duality map according to Eq. (6). While this direction was explored mathematically [26; 108; 52], it was not really implemented into the study of DFT yet.
  • The original motivation for devising the regularized KS scheme was to be able to prove this convergence. Yet, clearly, a guaranteed convergence is also beneficial for practical applications, so one could try and develop functional approximations $v_{\mathrm{xc},\varepsilon}(\rho)$ that are already adapted to the regularized KS iteration of Section V.1. A small step in this direction was presented in Section V.2 that demonstrated a way how to transform available exchange–correlation potentials such that they can be applied in the regularized KS scheme. But this must be seen just as a first approach to the question. Instead a construction of such approximations from first principles could also be tried. On the other hand, the practical implications of switching to the regularized form of the KS method have not really been explored yet.
  • Speaking of convergence, to date we do not have a convergence proof for the proximal-point iteration (Section IV.1) that is also the basis for the inverse KS method applied to periodic settings (Section VII). We can maybe expect it in the form of a fixed-point proof that is typical for Banach-space structured. We also note at this point, that the presented method can easily be adapted to other DFT settings, like spin DFT, current DFT, or QEDFT.
  • As mentioned, the homogeneous Sobolev spaces introduced in Section VI and their duals might have a central role in future reformulations of (regularized) DFT, but they are not that well studied in functional analysis, especially not on the dual side. What is known is that they are Hilbert spaces and thus Assumption 3 is secured. We noted in Section III.3 that they are already used in some practical schemes to enhance the Kohn–Sham method. Further functional theoretic insights on these spaces would thus be a valuable contribution to ongoing work. In the course of such study, the important questions of self-adjointness and boundedness below of the Hamiltonian and, in particular, the lower semicontinuity of the ensemble constrained-search density functional must also be clarified in order to have the correct setting for MY-regularized DFT according to Assumption 4.
  • The whole regularized KS framework in Section V.1 used the difference of the interacting functional $F^{1}_{\varepsilon}$ and the non-interacting $F^{0}_{\varepsilon}$, as typically done in KS-DFT. Yet, it was noted in Section VII that switching to a different guiding functional that already includes a priori information might be beneficial at least numerically. In a way we did this already in Section II.4 through the inclusion of the offset potential $v_{\mathrm{o}}$ that would include information about the external potential. But other, more complicated terms, like the Hartree mean-field energy $E_{\rm H}(\rho)$ could be included as well before regularization, thus choosing $(F^{0}+E_{\rm H})_{\varepsilon}$ as a regularized guiding functional. One only has to make sure that the functional is still convex and lower semicontinuous before regularization, in order to stay within Assumption 4. Further studies about this questions are already underway.
  • In the periodic setting, the computation of proximal points, respectively the inversion of Kohn–Sham densities based on MY techniques has recently become feasible for practical systems, see Section IV. However, the currently considered algorithm still suffers from a number of restrictions: (i) it is only applicable to insulating systems, (ii) proper preconditioner have not yet been developed and (iii) the choice of regularization parameters $\varepsilon$ requires careful manual tuning. Overcoming these restrictions is crucial to broaden the scope of MY-based numerical experiments for mathematical research on DFT as well as for practical KS inversion computations.
  • While recasting KS inversion via MY regularization yields an a priori orbital-free density-functional formulation, see Section IV, practical implementations of MY-based KS inversion (MYiKS) still rely on an explicit orbital parametrization to solve the proximal-point optimization problem Eq. (130). The same is true for many standard KS inversion algorithms [2; 109; 67]. Consequently, it would be advantageous to develop a mathematical framework in which the regularization is expressed directly in terms of the orbitals to provide a closer match between MY-based theory and practical inversion schemes. We remark that such an orbital-based formulation is also more naturally suited to provide a fully consistent treatment of the nonlocal term present in most pseudopotential approximations (see Section VII).
  • As previously discussed in Section VII, the close match between the mathematical framework and the numerical implementation in the periodic setting is remarkable. This enables joint analytical and numerical investigations where mathematical developments can be guided and tested using numerical experiments, and suggested theoretical relationships can be directly exploited in the development of inversion algorithms with improved efficiency and reliability. Considering the many possibilities for formulating MY-regularized DFT (e.g., the considerable flexibility in selecting appropriate function spaces), we further expect numerical investigations to be decisive for suggesting which theoretical extensions are worthwhile pursuing. Moreover, we remark that since the MY-based Kohn–Sham inversion framework can be viewed as a unified framework encompassing a wide range of inversion schemes, see Section IV, we expect the resulting mathematical and algorithmic advances from such studies to be broadly applicable and to directly contribute to the development of reliable and computationally efficient inversion schemes.
Data availability
No new data were created or analysed in support of this research.
Acknowledgements.
The authors thank Erik I. Tellgren, Mihály A. Csirik, Thiago Carvalho Corso, and Michael Ruggenthaler for valuable discussions. MP acknowledges support from the German Research Foundation (Grant SCHI 1476/1-1). MFH acknowledges support by the Swiss National Science Foundation (SNSF, Grant Nos. 221186 and 10002757) as well as the NCCR MARVEL, a National Centre of Competence in Research, funded by the SNSF (Grant No. 205602). AL and MP have received funding from the ERC-2021-STG under grant agreement No. 101041487 REGAL. AL and TH were funded by the Research Council of Norway through CoE Hylleraas Centre for Quantum Molecular Sciences Grant No. 262695.